### v3

# Unveiling the Influence of Amplifying Language-Specific Neurons

Abstract
Language-specific neurons, units in LLMs that strongly correlate with individual languages have been shown to influence model behavior by deactivating them. However, their role in amplification remains underexplored. This work investigates the effect of amplifying language-specific neurons through interventions across 18 languages, including low-resource ones, using three models primarily trained in different languages. We compare amplification factors by their effectiveness in steering to the target language using a proposed Language Steering Shift (LSS) evaluation score, then evaluate it on downstream tasks: commonsense reasoning (XCOPA, XWinograd), knowledge (Include), and translation (FLORES). The optimal amplification steering factors effectively steer output toward nearly all tested languages. Intervention using this factor on downstream tasks improves self-language performance in some cases but generally degrades cross-language results. These findings highlight the effect of language-specific neurons in multilingual behavior, where amplification can be beneficial especially for low-resource languages, but provides limited advantage for cross-lingual transfer.  [^1] The code and dataset are made available at https://github.com/tauimbz/lang-task-neuron.
Unveiling the Influence of Amplifying Language-Specific Neurons
Inaya Rahmanisa1, Lyzander Marciano Andrylie1, Mahardika Krisna Ihsani2, Alfan Farizki Wicaksono1, Haryo Akbarianto Wibowo2, Alham Fikri Aji2 1Faculty of Computer Science, Universitas Indonesia 2Department of Natural Language Processing, MBZUAI {inaya.rahmanisa, lyzander.marciano}@ui.ac.id, alfan@cs.ui.ac.id {mahardika.ihsani,haryo.wibowo,alham.fikri}@mbzuai.ac.ae
1 Introduction
Figure 1: Illustration of our methodology. There are three sequential stages involving: 1). Identification of language-specific neurons, 2). Finding the optimal amplifying steering factor, and 3). Evaluation of the optimal amplifying steering factor on downstream tasks to understand its impact on the models’ behavior.
Multilingual large language models (LLMs) have shown remarkable capabilities. However, they still continue to exhibit inconsistencies in downstream tasks (i.e., often producing biased or inaccurate outputs across different linguistic contexts), especially involving low-resource languages Dang et al. (2024). Understanding the root cause of these inconsistencies remains challenging due to limited transparency in how LLMs represent language-specific features and perform cross-lingual transfer.
In elucidating this issue, a few recent studies have begun to interpret model inner workings in multilingual contexts, revealing neuron-level correlations with language features and suggesting the existence of language-specific neurons Kojima et al. (2024); Tang et al. (2024); Zhao et al. (2024). These neurons have been shown to affect multilingual capability following deactivation and signs of language-steering capabilities in generation. Interventions are typically done to steer outputs by replacing the neuron values with fixed activation values, called steering factor, to control behavior. Most prior works use zero as the steering factor to deactivate targeted neurons. However, we argue that amplifying these neurons can offer deeper insight into their functional capabilities, particularly regarding performance improvements or task and language transfer.
Building on the success of cross-lingual transfer methods such as training and fine tuning Conneau et al. (2020); Kim et al. (2019), we aim to investigate whether neuron overlap reflects shared linguistic representations. To this end, we extend existing studies that have been largely focusing on self-language interventions Mondal et al. (2025); Wang et al. (2024) which showed limited gains, with cross-language interventions. We examine cross-lingual neuron behavior in closely related language pairs, distant language pairs, and low resource languages, to further explore whether neuron-level interventions can similarly provide insights on cross-lingual transfer, particularly revealing where and why such interventions succeed or fail.
This work is conducted by identifying and analyzing two types of neurons: language-specific neurons and language-activated (less-specific) neurons. We then compare various steering factors derived from different settings to evaluate their effect on the steerability of these neurons and analyze their impact on downstream task performance using neurons specific to the target language (self-intervention) and to every other languages (cross-intervention). We evaluate 18 languages including low-resource ones like Quechua, Haitian Creole, and Swahili, across models primarily trained in English and Chinese (Qwen), English (Gemma), and Southeast Asian languages (SeaLLMs) using XCOPA, XWinograd, and Include-lite for reasoning and knowledge tasks, and FLORES-200 for translation.
Several significant findings from this study are as follows:
  • Comparison between different steering factors using the proposed Language Steering Shift (LSS) reveals that patched steering factors are the most effective and statistically outperform test-time intervention factors. Amplifying language-specific neurons with these factors consistently directs outputs toward all target languages, achieving over 90% average success in language shifting.
  • Optimal amplification of language-specific neurons in self-intervention can reduce perplexity and improve task performance for certain languages, with gains reaching up to $1\%$ in some models, showing improved results for low-resource languages compared to Mondal et al. (2025).
  • Cross-lingual amplification can slightly reduce perplexity and enhance performance between closely related languages, but it generally leads to performance degradation in others, indicating limited effectiveness in cross-lingual transfer.
2 Related Work
Recent studies have explored neuron activations in language Kojima et al. (2024); Tang et al. (2024); Zhao et al. (2024), task Leng and Xiong (2025); Song et al. (2024), and mixed contexts Wang et al. (2024), often by inhibiting neurons to assess their roles. Kojima et al. (2024) showed that intervening on language-specific neurons can steer output language. Common intervention methods include activation patching Zhang and Nanda (2024), model editing Meng et al. (2023), and suppression via zeroing or percentile thresholds Mondal et al. (2025). While Mondal et al. (2025) has explored self-language intervention on downstream tasks, to the best of our knowledge, no study has yet focused on comparing amplification factors and provided improving insights into self- and cross-lingual interventions in downstream tasks.
3 Methodology
This section describes our method for identifying language-specific and language-activated neurons, introduces the Language Steering Shift (LSS) score to quantify language steering, and details our approach for determining the steering factors.
3.1 Neuron Identification
As illustrated in Figure 1, initially, we define neurons as the outputs of the activation function within the FFN modules in LLMs as they have been found to store knowledge or features Dai et al. (2022); Tang et al. (2024).
Identifying Language-Activated Neurons.
As a baseline method, we adapt a method from Wang et al. (2024) to define language-activated neurons, i.e., neurons whose activation value exceeds 0 throughout every sentence $s$ in the dataset of language $k$, denoted $S_{k}$, allowing a single neuron to be language-activated for multiple languages. The set of language-activated neurons in layer $i$ for language $k$, denoted by $A_{k}(i)$, is described by:
$A_{k}(i)=\left\{j\;\middle|\;\texttt{act}(i,j;s)>0,\;\;\forall s\in S_{k}\right\}\,.$
For neuron $j$, $\texttt{act}(i,j;s)$ is the activation value of the $j$-th neuron in the $i$-th layer when processing sentence $s$. These neurons are not further filtered, resulting in relatively large number of neurons and overlapping activations across languages. We refer to these as Baseline neurons.
Language-Specific Neurons.
language activation probablity entropy (LAPE) method (Tang et al., 2024) is adopted to identify language-specific neurons. The probability of activation for the $j$-th neuron within the $i$-th layer when processing language $k$, denoted by $p^{k}_{i,j}$, is estimated according to:
$\hat{p}^{k}_{i,j}=\mathbb{E}\left[\mathbb{I}(\texttt{act}({i,j})>0)\mid\text{language }k\text{ tokens}\right]\,.$
The expectation is taken over all tokens from sentences in language $k$, where $\mathbb{I}(.)$ is an indicator function that returns $1$ if the condition is true and 0 otherwise; and $\texttt{act}(i,j)$ denotes the activation value of the $j$-th neuron in the $i$-th layer. Let $\mathbf{p}_{i,j}=[p^{1}_{i,j},\dots,p^{k}_{i,j},\dots,p^{|L|}_{i,j}$] represent a list of activation probabilities for a set of languages $L$. These probabilities are then normalized to yield the L1-norm version, denoted by $p^{\prime}_{i,j}$. Using this normalized version, the LAPE score of the $j$-th neuron in the $i$-th layer is formally defined as:
$\text{LAPE}_{i,j}=-\sum_{k=1}^{|L|}p^{\prime k}_{i,j}\log(p^{\prime k}_{i,j})\,.$
We filter neurons whose language activation probability exceeds the $m$-th percentile. Finally, we select neurons that fall within the bottom $n\%$ of LAPE scores.
3.2 Language Steering Shift (LSS)
We propose Language Steering Shift (LSS) score to quantify the extent to which same-meaning answers are steered to the target language. Let $L$ be the set of languages used. For every language $k\in L$, we evaluate the dataset for language $k$ by performing: 1) a no intervention scenario, and 2) an intervention scenario where neurons corresponding to language $l$ are intervened where $l\in L\setminus\{k\}$. The LSS score for evaluating language $k$ when intervened upon by neurons specific to language $l$, denoted by $\texttt{LSS}(k,l)$, is calculated as follows:
$\texttt{LSS}(k,l)$ $=\mathbb{E}\left[\mathbb{I}\left(\delta(k,l)|_{\texttt{int}}-\delta(k,l)|_{\texttt{non}}>0\right)\right]\,$,
$\delta(k,l)$ $=\log p(\texttt{ans}_{l})-\log p(\texttt{ans}_{k})\,$,
where $\log p(\texttt{ans}_{x})$ denotes the log-probability assigned by the model to the answer in language $x$; the symbols $|_{\texttt{non}}$ and $|_{\texttt{int}}$ represent the no intervention and intervention scenario, respectively; $\mathbb{E}[.]$ is the expectation value across all sentences in the dataset. Intuitively, this metric measures the tendency of a model to produce the correct answer in target language $k$ given an incomplete sentence written in source language $l$ after intervening neurons corresponding to language $k$. For example, given the English sentence “Dublin is the capital city of Ireland”, the LSS score measures how much it shifts toward the Chinese translation, “Dublin is the capital city of 爱尔兰共和国”, when intervened using zh neurons. The underlined words represent the answers being analyzed.
3.3 Determining steering factors
We steer language-specific neurons in each layer with amplifying steering factors, which were determined in the identification stage (patched) and test-time intervention stage.
Patched Steering Factors. Let $v_{j,k}$ be the variable that contains the activation value of neuron $j$ that is specific to language $k$. In the identification stage involving the multilingual FLORES-200 dataset, the steering factors are computed from the max and median of activation values across all sentences within the dataset $S_{k}$ for language $k$. As a result, a steering factor is specific to neuron $j$ and language $k$, to which we patch the language-specific neurons. This process is formally described as:
$v_{j,k}\leftarrow\texttt{agg}\{s\in S_{k}|a_{j,s}\}\,$,
where $\texttt{agg}\in\{\texttt{max},\texttt{median}\}$; and $a_{j,s}$ represents the activation value of neuron $j$ averaged across all tokens in sentence $s$ during the identification stage. We follow Kojima et al. (2024) in using the median and extend their approach by incorporating a stronger aggregation function, max. We refer to these steering factors as patch max (pmax) and median (pmedian).
Test-Time Intervention Steering Factors. Let $v_{j,k,m}$ be the variable containing the activation of neuron $j$ for language $k$ when processing sentence $m$, i.e., a sentence from the test dataset. We define the steering value as the maximum over activation values within the sentence. Using this value, we apply two operations to the language-specific neurons: addition (+max) to assign with a value surpassing its activation range, and replacement (=max) to assign the maximum activation. This computation is formally represented as:
$\texttt{steer}_{j,m}\leftarrow\texttt{max}\{t\in T(m)|a_{j,m,t}\}\,$,
$\textbf{{=max}}\textrm{ : }v_{j,k,m}\leftarrow\texttt{steer}_{j,m}\,$,
$\textbf{{+max}}\textrm{ : }v_{j,k,m}\leftarrow v_{j,k,m}+\texttt{steer}_{j,m}\,$,
where $\texttt{steer}_{j,m}$ is the steering factor for neuron $j$ during inference on sentence $m$; $T(m)$ is the set of tokens in sentence $m$; and $a_{j,m,t}$ is the activation value of neuron $j$ associated with token $t$ on sentence $m$. As a comparison to our amplifying factors, we also use deactivating factors 0 and $10p$ (10th percentile) as a replacement of max, which has been similarly demonstrated by Mondal et al. (2025) to deactivate neurons. We refer to these as =0 and =10p, respectively.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
en 99.46 98.92 96.22 96.22 84.32 33.51 100.0 100.0 100.0 100.0 100.0 09.73 100.0 100.0 94.59 94.05 98.92 05.41
nl 100.0 100.0 92.97 90.81 67.57 32.97 100.0 100.0 100.0 100.0 68.65 15.14 100.0 100.0 93.51 89.73 61.08 16.22
ms 47.57 55.68 30.81 31.89 37.84 38.92 56.76 55.68 042.7 41.08 45.95 47.03 85.95 52.97 39.46 43.78 47.03 44.32
vi 100.0 100.0 080.0 75.14 23.78 18.92 100.0 100.0 96.76 96.22 14.05 10.81 100.0 100.0 87.57 76.76 14.59 08.65
jp 100.0 100.0 75.14 71.35 28.11 03.24 100.0 100.0 60.54 060.0 020.0 06.49 100.0 100.0 43.78 44.86 16.22 01.62
zh 100.0 100.0 48.11 43.78 14.05 05.41 100.0 100.0 87.03 81.08 33.51 09.73 100.0 100.0 56.76 56.76 49.19 01.08
fr 100.0 100.0 96.22 95.14 84.86 07.57 100.0 100.0 97.84 97.84 060.0 14.59 100.0 100.0 98.92 95.14 85.95 017.3
pt 99.46 100.0 96.76 94.05 72.43 09.73 100.0 100.0 98.92 98.92 077.3 09.19 100.0 100.0 96.76 91.35 082.7 04.32
ru 100.0 100.0 90.27 79.46 02.7 00.54 100.0 100.0 95.14 94.59 18.92 01.62 100.0 100.0 86.49 71.35 03.24 00.0
et 94.05 94.05 77.84 76.22 67.03 49.73 100.0 88.65 97.84 94.59 55.68 16.22 99.46 98.92 89.19 89.73 39.46 18.92
it 100.0 100.0 92.43 92.43 082.7 39.46 100.0 100.0 99.46 99.46 89.19 24.32 100.0 99.46 98.92 91.89 83.78 14.05
ta 100.0 100.0 01.08 01.08 00.0 00.0 100.0 98.38 50.81 45.95 00.0 00.0 100.0 100.0 01.08 01.08 00.0 00.0
th 100.0 100.0 41.62 41.62 00.54 00.54 100.0 100.0 65.41 64.32 01.08 00.0 100.0 100.0 25.41 022.7 00.54 00.0
tr 98.92 100.0 56.76 56.76 45.41 29.73 91.89 83.24 96.22 94.59 48.65 040.0 100.0 100.0 52.43 52.43 31.89 14.05
avg 89.30 89.91 65.08 63.06 40.76 18.02 89.91 88.40 79.24 77.91 42.20 13.66 92.36 90.09 64.32 61.44 40.97 9.73
Table 1: Language Steering Shift (LSS) scores from id to languages listed for amplifying (pmax, pmedian, =max, +max) and deactivating (=0, =10p) LAPE neurons across models. Bolded texts denote the highest values within a model.
Figure 2: Comparison of averaged LSS scores of various steering factors across languages using LAPE and Baseline neurons.
4 Experimental Settings
This section discusses the experimental settings used throughout the methodology. It consists of three core stages. First, the identification of neurons specifically activated by language and those exhibiting language-specific responses. Second, the process of finding the optimal amplifying steering factor for influencing these neural activities. Finally, the evaluation stage to determine the impact of this optimal amplifying steering factor on various downstream tasks to assess its efficacy.
Identification of Language-Activated Neurons and Language-Specific Neurons.
We identify language-activated (Baseline) and language-specific (LAPE) neurons by feeding the LLMs sentences from the dev split of FLORES-200 dataset Team et al. (2022) containing 30,00030{,}000 word-tokens per language. We assess 18 languages to ensure broader coverage of language specificity: en, nl, id, ms, vi, jp, zh, fr, pt, ru, et, ht, it, qu, sw, ta, th, and tr in small and large models, namely Qwen2.5 Instruct (0.5B and 7B), Gemma2 Instruct(2B and 9B), and SeaLLMs3 Chat (1.5B and 7B), totaling in 6 models (see Appendix B for details). For identifying LAPE neurons, we use a language activation probability threshold ($m$) set at the 95th percentile and a LAPE score threshold ($n$) of 1%. All hyperparameter values are adopted from Tang et al. (2024) as they have been demonstrated to effectively capture language-specific neurons.
Finding the Optimal Steering Factor.
We assess the determined steering factors on LSS score for the small models using a reconstructed version of MLAMA Kassner et al. (2021), a multilingual fact-probing dataset. MLAMA is reconstructed such that each language question includes answers from all other languages and is filtered to ensure that no answer in different languages are the same. For example, if both Indonesian and Malaysian answers are “Jerman” (i.e., Germany in English), that row is discarded. The dataset contains 185 instances per language, with 15 of the 18 available languages are used due to limitation in the dataset.
Evaluating the Optimal Steering Factor on Downstream Tasks.
We evaluate the model’s confidence and downstream task performances by conducting interventions using the optimal amplifying steering factors. We provide the deltas between the intervened and non-intervened performances to quantify the extent to which the interventions shift task outcomes. Interventions are performed in every layer of the model, with layer-wise amplification result and analysis provided in Appendix A.1.
We evaluate on commonsense reasoning task, which is considered one of the most difficult tasks for LLMs Li et al. (2022), using XCOPA Edoardo M. Ponti and Korhonen (2020) and XWinograd Muennighoff et al. (2022) dataset. For factual knowledge, we use Include-lite Romanou et al. (2024) dataset. Building on Mondal et al. (2025), who focused only on self-language intervention, we extend the analysis of these task performances to cross-lingual intervention. The devtest split of FLORES-200 is used to measure perplexity and BLEU for translation from English (en) to every other languages.
Translation tasks are performed using 2 types of prompts: targeted and non-targeted. Targeted prompts explicitly mention the target language, whereas non-targeted prompts do not. For example, when translating to id, the targeted prompt begins with "Translate from English to Indonesian.", whereas the non-targeted prompt begins with "Translate from English into the target language."
5 Result and Analysis
This section presents the findings and analysis of the optimal amplifying steering factors and their evaluation on model confidence and downstream tasks.
5.1 LSS Scores and Optimal Amplifying Steering Factor
Result of the Language Steering Shift (LSS) score shows that LAPE neuron amplification yields high scores for many languages, often reaching 100%, as shown in Table 1 for the intervention on id. These strong results are consistent across other languages, as detailed in Appendix E, indicating successful language shifting. However, interventions involving certain language pairs may result in lower LSS scores. For example, translation between id and ms in either direction tends to produce scores capped at 50% in most models. This is likely due to the high similarity between the two languages, which makes it more difficult for the models to distinguish between them. Notably, en remains easy to steer to, suggesting a strong influence that is possibly due to its dominance and widespread presence in LLM training data.
The optimal amplifying factors, defined as those that result in the highest aggregated LSS scores, are found to be pmax and pmedian. A one-way ANOVA followed by Tukey’s HSD test for post-hoc pairwise comparisons reveals that the patched steering factors pmax (mean = 0.9204) and pmedian ($\texttt{mean}=0.9095$) significantly outperform the other factors ($p<0.001$); Further results are provided in Appendix E. However, the difference between the two is not statistically significant ($p=0.98$), indicating that both are equally effective as optimal amplifying factors.
Test-time intervention factors, +max and =max, perform well mainly for high-resource languages such as en, pt, it, and fr, but show limited effectiveness for other languages, with the lowest performance observed for ta in Qwen and Gemma despite its substantial neuron presence. The deactivating factor =0 shows variation in results, with successful steering in some languages (e.g., en, nl, and it), while =10p demonstrates stronger deactivation capability, often resulting in LSS scores close to $0\%$ across many languages, align with findings from Mondal et al. (2025).
Compared to Baseline neurons, LAPE neurons yield higher average LSS scores across languages as illustrated in Figure 2. This suggests that LAPE is more effective at steering language, likely due to its stronger language specificity. In contrast, Baseline neurons show greater variability (see Appendix E); due to their lower specificity, deactivating these neurons with =10 and =10p sometimes result in higher LSS scores than amplifying them. This contrast further supports LAPE neurons as more reliable and consistent for steering towards corresponding language.
Model Include-lite XWinograd XCOPA
D O D O D O
S-7B $\mathbf{0.38}$ 0.11-0.11 $\mathbf{\phantom{-}1.27}$ 6.71-6.71 $\mathbf{-0.44}$ 5.49-5.49
S-1.5B $\mathbf{1.02}$ $\phantom{-}0.54$ $\mathbf{-0.76}$ 6.83-6.83 $\mathbf{-0.18}$ 3.18-3.18
Q-7B $\mathbf{0.26}$ 0.17-0.17 $\mathbf{-0.03}$ 7.19-7.19 $\mathbf{-0.22}$ 5.98-5.98
Q-0.5B $\mathbf{0.16}$ 0.36-0.36 $\mathbf{\phantom{-}0.17}$ 3.38-3.38 $\mathbf{-0.04}$ 1.69-1.69
G-9B $\mathbf{0.87}$ 0.77-0.77 $\mathbf{\phantom{-}0.29}$ 5.46-5.46 $\mathbf{\phantom{-}0.13}$ 2.24-2.24
G-2B $\mathbf{0.25}$ 1.03-1.03 $\mathbf{\phantom{-}0.15}$ 3.38-3.38 $\mathbf{\phantom{-}0.07}$ 1.43-1.43
Table 2: Average delta task accuracies after pmax intervention on LAPE neurons for self-intervention/diagonal (D) and cross-intervention/off-diagonal (O) neuron steering for every models used. Negative and positive values respectively represent decreased and increased accuracy from the baseline (non-intervention scenario). S denotes SeaLLMs3, Q = Qwen2.5 and G = Gemma2.
5.2 Impact of Neuron Amplification
Figure 3: Delta Include-lite accuracy under LAPE neurons on Gemma2 9B. Blue highlights reductions, whereas red highlights increases.
To further evaluate the optimal steering factors, we analyze their effectiveness in steering LAPE neurons using the most amplifying factor, pmax, across multiple downstream tasks and language modeling performance via perplexity. The next best factor, pmedian, shows comparable but less prominent results, while Baseline neuron amplification generally degrades performance in both self- and cross-intervention settings (see Appendix B for further results).
Figure 4: Delta XWinograd accuracy on Qwen2.5 7B (left) and Qwen2.5 0.5B (right) after amplification of LAPE neurons. Blue highlights reductions, whereas red highlights increases.
5.2.1 Knowledge and Reasoning Tasks
Overall performance of knowledge and reasoning tasks intervention as presented in Table 2 shows that pmax self-intervention can improve accuracy up to 1%, with consistent gains observed in Include-lite. XCOPA and XWinograd self-intervention may show less consistent improvements with only modest changes in some models. In contrast, cross-language interventions generally lead to performance drops on average. These results suggest that amplifying LAPE neurons in self-language on knowledge and reasoning tasks is potentially beneficial for enhancing performance, though their effectiveness in cross-lingual transfer appears limited.
Include-lite Evaluation.
The result in Figure 3 reveals that self-intervention improves performance or modestly changes. As a comparison, cross-intervention shows stronger degradation, visibly captured in et. Some models capture increasing accuracy across similar languages such as id and ms, though the correlations are not always clear as they often show better performance than self-intervention.
XWinograd Evaluation.
Figure 4 shows that self-intervention tends to be non-destructive, with some languages such as jp and pt, showing slight improvements across most models. However, the improved languages vary by models; for example, zh improves in Qwen2.5 0.5B but not in the 7B variant. Cross-lingual interventions degrade performance on average, particularly in larger models. Nevertheless, occasional improvements are observed in less related language pairs, such as zh when intervened by ru in Qwen2.5 0.5B (see Appendix H for additional results). A consistent trend across models is that amplifying en neurons reliably reduces accuracy, suggesting they may play a role in reasoning rather than language transfer.
Figure 5: Delta XCOPA accuracy after intevention of LAPE neurons for SeaLLMs3 7B (right). Row i depicts the initial language and column j is the intervention language.
XCOPA Evaluation.
Figure 5 shows that amplifying LAPE neurons does not consistently improve in self-intervention, but changes are generally reserved, indicating non-destructive effects. In contrast, cross-language amplification often leads to larger drops, visibly captured on the intervention of it, id, th, tr, and vi neurons, degrading other languages more than their own. Low-resource languages like qu, et, sw, and ta show self-intervention gains, particularly in SeaLLMs3. SeaLLMs3 and Qwen2.5, having shared model architecture Zhang et al. (2024), display very similar patterns, suggesting architecture may contribute to neuron behavior. Overall, XCOPA deltas reveal clearer signs of degradation than improvement.
Lang Targeted Prompt Non-targeted Prompt
G-2B G-9B Q-0.5B Q-7B S-1.5B S-7B G-2B G-9B Q-0.5B Q-7B S-1.5B S-7B
et -0.44 (5.27) ++0.04 (5.06) -0.04 (1.24) -0.74 (4.09) -0.15 (2.32) -0.81 (3.76) ++0.07 (0.45) -0.01 (0.24) ++0.40 (0.49) ++1.78 (0.76) ++1.01 (0.87) ++2.56 (0.43)
fr -1.23 (26.49) -4.19 (41.90) ++0.08 (13.93) -0.16 (27.24) ++0.21 (18.23) -1.36 (30.47) ++8.32 (0.76) ++3.97 (0.57) ++12.39 (0.52) ++14.40 (13.75) ++18.38 (1.43) ++24.69 (0.32)
ht -0.94 (2.53) -1.06 (2.21) -0.08 (1.55) -0.15 (3.35) -0.19 (3.34) -2.03 (5.12) ++0.01 (0.17) -0.01 (0.13) ++0.78 (0.25) ++2.19 (0.35) ++1.53 (0.56) ++2.53 (0.43)
id -0.53 (18.96) -1.60 (17.89) ++0.49 (7.34) ++2.15 (21.29) -0.64 (20.46) ++0.61 (30.02) ++1.09 (0.55) ++0.21 (0.29) ++5.72 (0.65) ++23.88 (1.13) ++14.07 (1.92) ++26.36 (0.52)
it -0.67 (15.79) -2.44 (25.30) ++0.03 (6.34) -1.23 (15.90) -1.04 (8.23) -0.49 (17.42) ++3.66 (0.48) ++0.20 (0.25) ++1.74 (0.63) ++13.44 (1.16) ++9.62 (1.15) ++15.55 (0.44)
jp ++0.01 (0.11) -0.38 (0.46) ++0.15 (0.05) ++0.20 (0.05) ++0.36 (0.14) ++0.06 (0.44) ++0.09 (0.02) -0.02 (0.03) ++0.09 (0.06) ++0.06 (0.09) ++0.24 (0.11) ++0.06 (0.08)
ms -0.38 (13.51) -0.89 (35.14) ++0.08 (4.85) -0.24 (16.13) -0.25 (15.39) -0.85 (24.82) ++0.69 (0.49) ++0.08 (0.31) ++3.32 (0.61) -0.37 (1.01) ++9.05 (1.54) ++18.64 (0.50)
nl ++0.05 (13.08) -2.72 (21.08) ++0.15 (4.52) ++1.01 (14.30) ++0.33 (8.57) -0.66 (14.67) ++4.49 (0.45) ++0.06 (0.24) ++2.31 (0.55) ++9.84 (0.85) ++5.36 (1.20) ++11.01 (0.41)
pt -2.42 (25.08) -1.77 (44.80) ++0.41 (13.98) -1.02 (36.48) ++0.39 (12.19) ++0.23 (30.99) ++10.11 (0.61) ++12.45 (0.44) ++9.45 (0.61) ++23.96 (1.33) ++15.32 (1.47) ++18.47 (0.58)
qu -0.06 (0.34) ++0.39 (0.21) ++0.05 (0.46) -0.12 (1.39) -0.28 (1.05) -0.30 (0.64) ++0.12 (0.44) ++0.16 (0.34) ++0.30 (0.46) ++0.38 (0.75) -0.00 (0.93) ++0.11 (0.63)
ru -0.33 (13.72) ++1.66 (16.16) ++0.26 (5.67) -0.51 (16.88) -0.99 (8.90) -3.64 (17.02) -0.02 (0.38) ++1.61 (0.24) -0.93 (5.22) ++14.03 (2.20) ++9.62 (0.75) ++11.58 (4.77)
sw -0.34 (7.62) ++2.71 (12.15) -0.24 (0.58) -0.48 (2.39) -0.37 (1.56) -0.06 (3.06) ++0.03 (0.45) -0.05 (0.26) ++0.35 (0.51) ++1.01 (0.78) -0.06 (1.07) ++1.59 (0.36)
ta -0.66 (2.21) -0.16 (5.58) -0.01 (0.13) -0.05 (0.28) -0.05 (0.42) -0.14 (0.59) -0.01 (0.40) -0.10 (0.23) -0.17 (0.39) -0.23 (0.45) -0.21 (0.64) ++0.12 (0.29)
th -0.45 (8.38) -0.63 (19.93) -0.49 (2.19) -0.40 (8.08) -0.10 (3.69) -1.20 (6.61) ++0.06 (0.45) ++0.06 (0.28) ++0.71 (0.48) ++6.96 (0.77) ++2.36 (1.02) ++5.29 (0.43)
tr -2.49 (22.69) -5.95 (32.65) ++0.10 (10.71) ++4.73 (23.58) ++0.08 (18.12) ++0.89 (26.70) ++0.77 (0.33) ++4.42 (0.16) ++9.59 (0.52) ++25.26 (0.74) ++15.65 (1.65) ++21.40 (0.36)
vi -0.35 (0.53) -2.23 (2.89) ++0.10 (0.13) -0.53 (1.53) -0.17 (0.62) -0.30 (0.56) ++0.00 (0.09) ++0.03 (0.09) -0.02 (0.20) ++0.28 (0.19) -0.10 (0.32) ++0.95 (0.24)
Table 3: BLEU score changes of self-language pmax intervention from en to target languages with Targeted and Non-targeted prompts on the FLORES-200 dataset. Each values are changes with increases shown in green and decreases in red. The subscripted values to the right of each number indicate the BLEU score without intervention. Q denotes Qwen2.5, G denotes Gemma2 and S denotes SeaLLMs-v3.
Figure 6: BLEU deltas for non-targeted pmax intervention in Qwen2.5 0.5B. Row $i$ is the target language and column $j$ is the intervention language. For example, row $x$ col $y$ represents BLEU score when translating English to language $x$ intervened by $y$ neurons.Blue highlights reductions, whereas red highlights increases.
5.2.2 Translation Task
Table 3 shows that BLEU scores increase after amplifying LAPE neurons with pmax for translating English to 17 other languages with non-targeted prompt. The results suggest that amplifying language-specific neurons enhances the model’s ability to generate text in the target language, including low-resource languages such as qu, ht, and sw, across most models. In contrast, the improvements from targeted prompts are less pronounced. Smaller models like Qwen2.5 0.5B and SeaLLMs3 1.5B tend to exhibit more consistent gains, which may indicate a higher degree of neuron locality in these models. The difference in performance between targeted and non-targeted prompts may suggest that amplifying these neurons improves the generation of the target language, though not always with high accuracy for translation. In contrast to amplification of Baseline neurons, which often degrades BLEU scores (Appendix G), LAPE demonstrates the ability to improve performance by targeting language-specific neurons.
To support the increasing BLEU findings, particularly in non-targeted prompt, we compare the performance with the cross-intervention setting presented in Figure 6. Self-intervention results in increasing BLEU scores compared to cross-intervention. Notably, using ru as the intervention language reduces performance in the cross-intervention case, while other languages show more modest increase compared to self-intervention. The observed improvements across languages may relate to linguistic similarity. For example, some increases occur when translating to id with ms as the intervention language, and vice versa; or when translating to pt with fr or it, and to fr with ht.
5.2.3 Perplexity Changes
Self-intervention of LAPE neurons using pmax can modestly reduce perplexity in most languages, showing success in improving language modeling performance as illustrated in Figure 7. Perplexity tends to increase more in cross-intervention settings than in self-intervention. However, the extent of the increase is not always related to the similarity of the language pair; For example, intervention between highly similar and same-family languages like id and ms does not lower perplexity, while less-related pairs like fr on en can. Interestingly, smaller models exhibit larger increases in perplexity, a pattern consistent with findings from the translation task intervention. This may suggest that smaller models contain more localized neurons responsible for language identity (see Appendix F for additional results).
The positive effects of perplexity reduction from LAPE neuron amplification are further supported by a comparison with Baseline neuron amplification. Intervention using Baseline neurons shows a similar pattern of increased perplexity in cross-intervention settings, although the increases are more pronounced. In self-intervention, amplification with Baseline neurons barely reduces perplexity, indicating no improvement in language modeling performance. These results highlight the advantage of LAPE neurons, whose specificity contributes to their ability to enhance language modeling performance. Detailed results for Baseline neuron interventions are provided in Appendix F.
5.3 Overlap Patterns in Language-Specific Neurons
As illustrated in Figure 8, although LAPE neurons are inherently tailored to individual languages, overlaps still exist across languages, albeit at significantly lower proportions. Compared with Baseline neurons with higher overlap proportions (Appendix D), the stronger specificity of LAPE neurons significantly improve self-intervention and capture patterns in cross-intervention for perplexity and task performance, indicating that better specificity leads to better performance for neuron-level intervention.
Figure 7: PPL change with pmax steering factor under LAPE neurons on Gemma2 9B. Row $i$ is the input language and column $j$ is the language of intervention.
A strong overlap is captured in LAPE neurons across mutually intelligible languages like Indonesian (id) and Malaysian (ms), likely due to their linguistic similarity Ahmad and Pustaka (2011). Languages sharing writing systems, such as Chinese (zh) and Japanese (jp) Taylor et al. (1995), also show overlap, consistent with Kojima et al. (2024). Romance languages, i.e., Italian (it), French (fr), and Portuguese (pt), share neurons, and Vietnamese (vi) overlaps moderately with Thai (th) and slightly with id, ms, jp, and zh, likely due to regional or historical ties. Haitian Creole (ht) overlaps with French, reflecting shared roots Prou et al. (2012). Surprisingly, low-resource languages like Quechua (qu), Haitian Creole (ht), and Swahili (sw) also show overlap, possibly indicating more general or shared neurons.
Figure 8: Symmetric heatmaps of overlap LAPE neuron proportions in SeaLLMs3 7B using jaccard distance.
6 Conclusion
This study proposes the Language Steering Shift (LSS) metric to compare steering factors of language-specific (LAPE) neuron intervention and identify the optimal one, then evaluates its impact in downstream tasks and language modeling. Optimal amplification steers outputs toward target languages with over $90\%$ success on average. Self-language intervention modestly improves or maintains performance in language modeling measured by perplexity, translation, knowledge, and reasoning tasks. Cross-language intervention generally lowers performance, though gains can appear between closely related languages in perplexity and BLEU scores for translation. Compared to less specific language-activated neurons, LAPE neurons yield clearer, more consistent improvements, suggesting that greater neuron specificity enhances multilingual performance when intervened. Overall, these findings suggest that amplifying LAPE neurons in self-language can enhance a model’s language capabilities, although not always improving task accuracy. Furthermore, the consistent degradation from cross-language interventions highlights their limited effectiveness for cross-lingual transfer.
Limitations
This study is limited to decoder-only models in 0.5B to 9B parameters and limited sets of tasks. Methods to identify language-specific neurons are also limited to LAPE, there are still gaps on the effect of amplification on neurons obtained by fine-grained attribution methods such as Integrated Gradients.
Appendix A Ablation Study
A.1 Layer-Wise Amplification on LSS Score
We perform layer-wise amplification on Gemma2 2B model (26 layers) on LSS score using the reconstructed MLAMA dataset. Consistent throughout each language of intervention as captured by the average LSS scores across all languages in Figure 68, the highest average is found in the 19th layer, and the lowest is in the 10th layer.
Later layers are more likely to result in higher LSS score, especially from the 15th layer until just before the last layer. First few layers may also result in high LSS scores, though not as high. The middle layers result in relatively low LSS scores, often reaching 0 in most languages, with the exception of tr whose the lowest LSS is found in the 18th layer. We observe that some languages from the same family share similar layer-wise LSS scores, such as id, ms, vi and pt, it, fr.
This result aligns with findings from Tang et al. (2024) and Kojima et al. (2024) as language-specific neurons are mostly found in the first and last layers, excluding the very last one. Results for layer-wise amplification for each language are provided in Figure 69 (en), 71 (et), 72 (nl), 73 (id), 70 (ms), 74 (vi), 75 (jp), 76 (zh), 77 (fr), 78 (pt), 79 (ru), 80 (it), 81 (ta), 82 (th), 83 (tr).
A.2 Identifying Language-Specific Neurons from XWinograd to Evaluate XWinograd Performance Under Intervention
In this ablation study, we explore the behavior of amplification when the language-specific neurons are identified from the same dataset used for task evaluation. Instead of using FLORES-200 as in our original approach, we identify the language-specific neurons using the XWinograd dataset. Once identified, we amplify the neurons with pmax steering factor to evaluate the exact same XWinograd dataset from which the neurons were identified. The result in Figure 20 shows that there are no indication of improvement in the self-intervention nor cross-intervention in similar languages, leading to the conclusion that identifying and steering language-specific neurons from the same dataset may not better improve task performance in the respective languages. This result captures less visible patterns, which may be due to the fact that the XWinograd dataset contains significantly fewer tokens than FLORES-200. This scarcity makes LAPE less effective at identifying language-specific neurons, and consequently, it struggles to improve task performance.
Appendix B Model and Language Details
B.1 Models and Language Neurons Used
We use models trained in different primary languages: Qwen2.5 Instruct Qwen et al. (2025) 0.5B and 7B (primarily trained in Chinese and English), Gemma2 Instruct Team et al. (2024) 2B and 9B (primarily trained in English), SeaLLMs-v3 Chat Zhang et al. (2024) 1.5B and 7B (primarily trained in Southeast Asian languages). Qwen2.5 and SeaLLMs-v3 share the same model architecture, whereas Gemma2 has a different architecture. We employ neurons specific to 18 languages, English (en), Dutch (nl), Indonesian (id), Malay (ms), Vietnamese (vi), Japanese (jp), Chinese (zh), French (fr), Portuguese (pt), Russian (ru), Estonian (et), Haitian Creole (ht), Italian (it), Quechua (qu), Swahili (sw), Tamil (ta), Thai (th), and Turkish (tr). The selected languages account for different language domains associated with the models and represent both high and low-resource languages to further analyze their behavior.
B.2 Experimental Settings for Finding the Most Optimal Steering Factors
To find the most optimal steering factors, we reconstruct MLAMA dataset Kassner et al. (2021) using 15 languages due to limitation in the dataset. The 15 languages are English (en), Dutch (nl), Indonesian (id), Malay (ms), Vietnamese (vi), Japanese (jp), Chinese (zh), French (fr), Portuguese (pt), Russian (ru), Estonian (et), Italian (it), Tamil (ta), Thai (th), and Turkish (tr). In the reconstructed MLAMA dataset, each question has 15 answers in different languages, with no lexical overlap.
We evaluate the proposed Language Steering Shift (LSS) scores on the reconstructed MLAMA dataset to measure how much an answer is shifted to the target language after intervened by the steering factors. This evaluation is only performed on the small versions of each model (Qwen2.5 0.5B, Gemma2 2B, and SeaLLMs-v3 1.5B) as they have already resulted in very high LSS scores, showing strong capability in steering to target languages.
B.3 Experimental Settings for Task Evaluation
We evaluate every models (6 in total) mainly using the pmax steering factor on several tasks. For translation task and perplexity, we use the devtest split of the FLORES-200 dataset on the same 18 languages the neurons were identified from. The translation evaluation is performed from English to every other 17 languages as English is considered the dominant language due to its widespread use as a multilingual benchmarks. Translation tasks are performed using 2 types of prompt: targeted and non-targeted. Targeted prompts explicitly mention the target language, wherease non-targeted prompts do not. For example, for translation to id, the targeted prompt is as follows,
    "Translate from English to Indonesian.
    English: {question}
    Indonesian:"
whereas the non-targeted prompt is as follows,
    "Translate from English into the
    target language.
    English: {question}
    Target language:"
For commonsense reasoning task, we evaluate XCOPA Edoardo M. Ponti and Korhonen (2020) and XWinograd Muennighoff et al. (2022), each containing 2-option questions. XCOPA dataset aims to choose the more plausible cause or effect of a given premises with a focus on lower-resource languages including Estonian (et), Haitian Creole (ht), Indonesian (id), Italian (it), Quechua (qu), Swahili (sw), Tamil (ta), Thai (tah), Turkish (tr), Vietnamese (vi), and Chinese (zh). XWinograd aims to resolve an ambiguous pronoun using commonsense knowledge, available in English (en), French (fr), Japanese (jp), Portuguese (pt), Russian (ru), and Chinese (zh).
For knowledge and reasoning task, we evaluate Include-lite dataset Romanou et al. (2024), containing 4-option multiple-choice-questions (MCQ) extracted from academic and professional exams, covering 57 topics, including regional knowledge. The languages used for this dataset are limited to the available languages in the obtained language-specific neurons set, which are the following: Dutch (nl), Indonesian (id), Malay (ms), Vietnamese (vi), Japanese (jp), Chinese (zh), French (fr), Portuguese (pt), Russian (ru), Estonian (et), Italian (it), Tamil (ta), and Turkish (tr).
Appendix C Proportions of LAPE and Baseline neurons
Throughout all models, the proportions of LAPE neurons are presented in Table 27, Baseline neurons in Table 28. Proportions of LAPE neurons show a significant variability across languages and models. Some lower-resource languages (e.g. qu, sw, ta) show very high proportions for certain models, while others (e.g. en, nl, ms, vi) generally show lower proportions. This may suggest these languages need more dedicated neurons due to linguistic complexity or limited pre-training exposure. In contrast, en has the fewest neurons across models, consistent with Tang et al. (2024), likely because English’s dominance in pre-training leads to more distributed, broadly learned features.
Appendix D Neuron Overlaps
We calculate the Jaccard distance for neuron overlap between languages for LAPE neurons (Figure 10, 13, 14, 11, 12) and Baseline neurons (Figure 15, 18, 19, 16, 17).
Appendix E LSS Scores
The complete LSS scores for both Baseline and LAPE neurons are described in Table 5 to Table 18. The aggregated LSS scores for each steering factor is presented in Table 26. Further statistical testing to determine the significance of the optimal steering factor is provided in 19. Visualization of token log-probs for same-meaning answers in different languages are provided in both intervened and non-intervened LLMs in Figure 84 and 85.
Appendix F Perplexity Changes After Amplification
Perplexity changes for pmax intervention of LAPE neurons are provided in Figure 58, 59, 56, 57, 60, and 61; pmedian intervention are provided in Figure 64, 63, and 62. Perplexity changes for pmax intervention of Baseline neurons are provided in Figure 66, 65, and 67.
Appendix G BLEU Score Changes for Translation Tasks After Amplification
For the self-language intervention, BLEU score changes for pmax intervention of LAPE neurons with non-targeted prompt is provided in Table 21. Sample of BLEU delta heatmap (self- and cross-intervention) is provided in Figure 9. BLEU score changes for pmedian intervention of LAPE neurons with targeted prompt is provided in Table 20. pmax intervention on Baseline neurons with targeted prompt is provided in Table 22. Samples of translation output for both targeted and non-targeted prompts are presented in Figure 86 and 87.
Appendix H Delta Accuracies of Reasoning and Knowledge Tasks After Amplification
Table 23 reveals the baseline (no intervention) XWinograd accuracy. Delta accuracies of XWinograd dataset under LAPE neurons when steered with pmax are provided in Figure 21, 22 and 23; intervention results for pmedian are provided in Figure 24, 25, and 26. Delta XWinograd accuracies of pmax intervention on Baseline neurons are provided in Figure 29, 30, and 31.
Table 24 reveals the baseline (no intervention) XCOPA accuracy. Delta accuracies of XCOPA dataset under LAPE neurons when steered with pmax are shown in Figure 32, 33, 34, and 36; intervention results for pmedian are shown in Figure 39, 38, and 40. Delta XCOPA accuracies of pmax intervention on Baseline neurons are provided in Figure 41, 42, and 43.
Table 25 reveals the baseline (no intervention) Include-lite accuracy. Delta accuracies of Include-lite dataset under LAPE neurons when steered with pmax are shown in Figure 47, 46, 45, 44, 49, and 48, intervention results for pmedian are provided in Figure 51, 50, and 52. Delta Include-lite accuracies of pmax intervention on Baseline neurons are provided in Figure 54, 53, and 55.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 96.22 97.84 95.14 94.05 92.43 92.97 99.46 99.46 79.46 78.38 31.35 81.62 99.46 100.0 95.14 94.05 51.89 60.54
nl 90.81 88.11 80.54 81.08 71.89 78.92 71.89 062.7 42.16 44.32 31.35 58.92 94.05 97.3 76.22 77.84 042.7 43.78
ms 40.54 46.49 31.89 30.81 29.73 34.05 29.73 25.95 31.35 31.35 41.08 040.0 040.0 42.16 31.89 32.43 36.22 23.78
vi 89.73 90.81 76.22 76.22 54.59 59.46 81.08 72.97 31.35 29.73 09.19 14.59 98.38 100.0 72.43 70.81 12.97 16.22
jp 87.57 88.11 63.78 68.11 29.73 28.11 79.46 91.89 042.7 41.62 21.62 38.38 97.3 97.3 042.7 36.76 10.81 05.41
zh 87.03 95.68 54.05 55.68 48.11 32.97 78.92 90.27 27.03 26.49 27.03 32.97 100.0 100.0 25.95 20.54 13.51 02.16
fr 92.97 95.14 78.38 82.16 67.57 74.05 91.89 94.59 77.84 080.0 037.3 47.57 100.0 99.46 80.54 78.38 52.43 54.05
pt 91.89 89.19 75.14 74.05 61.08 69.19 95.14 98.38 78.92 81.08 35.14 060.0 88.65 88.65 71.35 73.51 27.03 32.43
ru 92.97 97.84 57.84 060.0 34.59 037.3 88.11 90.81 27.57 26.49 05.95 12.43 99.46 100.0 41.08 38.38 03.78 05.41
et 94.05 92.97 72.97 73.51 60.54 68.11 61.08 62.7 32.43 30.81 020.0 21.08 92.97 95.14 69.73 69.73 30.27 35.68
it 98.92 96.76 83.78 87.03 66.49 74.59 93.51 98.38 75.68 74.59 47.57 67.03 95.14 97.84 75.68 75.14 44.86 52.43
ta 73.51 68.11 01.08 01.08 00.54 00.0 10.81 05.95 06.49 04.32 00.0 00.0 80.54 76.76 00.54 00.54 00.0 00.0
th 95.14 95.68 14.05 14.59 07.57 04.86 52.97 60.54 11.35 11.89 00.54 00.54 99.46 100.0 05.41 08.11 01.62 01.08
tr 92.97 91.89 56.22 55.68 41.08 45.95 45.95 52.97 24.32 29.19 37.84 37.84 94.59 96.76 45.95 48.11 19.46 28.11
LAPE
en 99.46 98.92 96.22 96.22 84.32 33.51 100.0 100.0 100.0 100.0 100.0 09.73 100.0 100.0 94.59 94.05 98.92 05.41
nl 100.0 100.0 92.97 90.81 67.57 32.97 100.0 100.0 100.0 100.0 68.65 15.14 100.0 100.0 93.51 89.73 61.08 16.22
ms 47.57 55.68 30.81 31.89 37.84 38.92 56.76 55.68 42.70 41.08 45.95 47.03 85.95 52.97 39.46 43.78 47.03 44.32
vi 100.0 100.0 80.00 75.14 23.78 18.92 100.0 100.0 96.76 96.22 14.05 10.81 100.0 100.0 87.57 76.76 14.59 08.65
jp 100.0 100.0 75.14 71.35 28.11 03.24 100.0 100.0 60.54 60.0 20.0 06.49 100.0 100.0 43.78 44.86 16.22 01.62
zh 100.0 100.0 48.11 43.78 14.05 05.41 100.0 100.0 87.03 81.08 33.51 09.73 100.0 100.0 56.76 56.76 49.19 01.08
fr 100.0 100.0 96.22 95.14 84.86 07.57 100.0 100.0 97.84 97.84 60.00 14.59 100.0 100.0 98.92 95.14 85.95 17.30
pt 99.46 100.0 96.76 94.05 72.43 09.73 100.0 100.0 98.92 98.92 77.30 09.19 100.0 100.0 96.76 91.35 82.7 04.32
ru 100.0 100.0 90.27 79.46 02.70 00.54 100.0 100.0 95.14 94.59 18.92 01.62 100.0 100.0 86.49 71.35 03.24 00.00
et 94.05 94.05 77.84 76.22 67.03 49.73 100.0 88.65 97.84 94.59 55.68 16.22 99.46 98.92 89.19 89.73 39.46 18.92
it 100.0 100.0 92.43 92.43 82.70 39.46 100.0 100.0 99.46 99.46 89.19 24.32 100.0 99.46 98.92 91.89 83.78 14.05
ta 100.0 100.0 01.08 01.08 00.00 00.00 100.0 98.38 50.81 45.95 00.00 00.00 100.0 100.0 01.08 01.08 00.00 00.00
th 100.0 100.0 41.62 41.62 00.54 00.54 100.0 100.0 65.41 64.32 01.08 00.00 100.0 100.0 25.41 22.70 00.54 00.00
tr 98.92 100.0 56.76 56.76 45.41 29.73 91.89 83.24 96.22 94.59 48.65 40.00 100.0 100.0 52.43 52.43 31.89 14.05
Table 4: Language Steering Shift (LSS) scores from id to languages listed for amplifying (pmax, pmedian, =max, +max) and deactivating factors (=0, =10p) across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
nl 63.24 68.65 30.81 32.97 05.41 10.27 63.78 70.81 37.84 40.54 29.19 11.89 90.27 97.3 017.3 19.46 10.27 08.11
id 67.03 69.73 24.32 21.08 11.35 12.97 077.3 77.84 65.41 65.95 060.0 52.43 87.03 96.22 15.68 16.22 05.95 04.86
ms 47.57 47.57 24.86 23.78 08.65 06.49 54.59 042.7 28.11 28.65 28.65 24.86 77.84 87.03 12.43 07.03 05.41 03.78
vi 74.59 67.57 28.11 25.41 07.57 07.57 73.51 70.27 39.46 41.62 09.19 05.95 95.14 94.05 10.81 10.27 04.86 05.95
jp 34.59 45.95 07.03 06.49 01.08 01.08 57.84 66.49 29.19 28.11 27.03 07.03 84.86 97.3 00.54 00.54 01.62 00.54
zh 29.73 81.08 05.95 06.49 10.27 01.62 50.27 83.24 09.73 09.19 21.08 05.95 82.16 97.84 00.0 00.0 14.59 01.62
fr 67.57 71.89 47.03 46.49 38.92 37.84 85.41 88.65 58.92 58.92 56.22 22.16 88.11 98.92 037.3 35.14 21.62 18.92
pt 71.35 71.35 07.03 06.49 03.78 06.49 88.11 91.35 66.49 66.49 64.86 31.89 77.84 94.05 08.65 08.65 07.03 04.32
ru 58.38 70.27 04.32 04.86 02.7 03.78 78.38 83.24 020.0 21.08 04.86 04.32 92.97 97.84 09.19 05.41 01.62 01.08
et 94.59 94.59 17.84 13.51 12.97 10.81 58.38 62.16 30.27 23.78 05.41 03.78 96.22 96.22 06.49 07.03 06.49 02.16
it 74.59 75.14 30.27 29.73 23.24 020.0 89.19 91.35 62.16 61.62 58.38 38.38 87.57 95.14 18.92 21.62 12.43 03.24
ta 10.81 09.19 00.0 00.0 00.0 00.0 3.24 01.08 01.08 01.08 00.0 00.0 19.46 11.35 00.0 00.0 00.0 00.0
th 80.0 78.38 00.54 01.62 00.54 00.54 31.35 43.24 01.08 02.16 00.0 00.0 95.14 98.92 01.62 01.62 01.08 00.0
tr 71.35 69.73 18.92 16.76 03.78 07.03 35.68 35.14 08.11 06.49 16.76 09.73 83.78 98.38 04.86 07.03 04.32 03.24
LAPE
nl 97.84 96.76 41.08 40.54 53.51 21.62 99.46 96.76 95.14 94.59 63.24 10.27 99.46 100.0 72.97 70.81 54.05 11.89
id 98.92 99.46 28.11 23.78 81.62 14.59 98.92 94.59 95.14 92.97 71.89 28.11 100.0 100.0 52.97 56.76 060.0 03.78
ms 98.38 98.92 17.84 19.46 45.95 10.27 97.3 94.59 94.59 94.05 46.49 14.59 98.92 99.46 32.43 33.51 32.97 04.32
vi 100.0 100.0 48.11 38.38 18.38 03.24 98.92 98.38 76.76 79.46 017.3 09.73 100.0 100.0 71.35 44.86 15.14 02.16
jp 99.46 99.46 08.65 08.11 25.95 00.54 100.0 99.46 41.08 40.54 35.68 08.65 100.0 100.0 01.08 00.54 08.65 00.0
zh 99.46 99.46 07.03 00.0 12.97 07.03 95.68 99.46 60.54 61.08 38.38 06.49 100.0 100.0 10.81 02.16 23.24 04.86
fr 96.22 96.76 077.3 74.05 70.81 07.57 100.0 98.92 79.46 77.84 70.81 16.22 100.0 100.0 76.76 70.27 65.95 01.62
pt 99.46 99.46 057.3 40.54 062.7 03.78 100.0 100.0 80.54 81.62 72.97 26.49 100.0 100.0 74.05 61.62 62.16 03.78
ru 100.0 100.0 16.22 11.89 02.7 01.62 99.46 99.46 83.24 080.0 10.81 02.16 100.0 100.0 022.7 17.84 05.41 00.54
et 96.76 94.59 15.68 14.05 56.76 18.38 98.92 96.76 95.68 95.68 39.46 05.95 97.84 97.84 31.89 31.89 57.84 25.95
it 99.46 99.46 060.0 44.86 68.11 13.51 100.0 99.46 87.57 87.03 69.73 09.73 99.46 99.46 61.62 58.92 76.76 03.24
ta 99.46 99.46 00.0 00.0 00.0 00.0 97.3 68.11 11.89 11.35 00.0 00.0 99.46 99.46 00.0 00.0 00.0 00.0
th 100.0 100.0 03.24 02.7 00.0 00.0 100.0 95.68 34.59 32.97 00.0 00.0 100.0 100.0 01.62 01.08 01.08 00.0
tr 98.92 98.92 15.14 13.51 34.59 09.73 86.49 80.54 72.97 67.57 42.16 13.51 97.84 97.84 19.46 19.46 32.97 03.24
Table 5: Language Steering Shift (LSS) scores from en to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 100.0 99.46 95.68 92.97 78.92 88.65 99.46 99.46 100.0 100.0 84.86 96.76 100.0 93.51 99.46 97.84 78.38 79.46
nl 69.19 65.95 66.49 69.19 65.41 70.81 91.35 91.35 78.92 84.32 74.59 81.62 83.78 71.35 74.59 74.05 58.38 69.19
id 63.78 64.86 75.68 62.16 060.0 67.57 96.76 95.68 94.05 94.59 84.32 92.97 70.81 67.57 71.89 72.97 59.46 63.78
ms 52.43 49.19 38.38 34.59 31.89 41.62 84.32 84.32 69.19 69.19 66.49 84.86 45.95 50.27 37.84 37.84 25.95 27.03
vi 89.73 89.73 63.24 66.49 37.84 54.59 98.92 98.38 86.49 86.49 30.81 71.89 99.46 98.92 76.76 080.0 27.03 29.19
jp 85.95 89.19 50.27 49.73 14.59 017.3 95.14 95.14 78.92 79.46 29.19 49.73 99.46 99.46 040.0 35.68 02.16 02.16
zh 83.78 96.76 042.7 52.97 41.62 27.03 97.3 97.3 78.92 81.62 28.65 45.95 98.92 100.0 27.03 25.95 24.32 02.7
fr 89.19 92.43 077.3 74.05 75.14 74.59 99.46 98.38 94.05 94.05 66.49 92.97 94.05 98.38 83.24 082.7 61.62 66.49
pt 77.84 83.78 64.32 65.41 58.92 65.41 98.38 98.38 97.84 97.84 76.22 88.11 84.86 91.89 69.73 71.35 41.08 48.11
ru 097.3 98.38 37.84 45.41 022.7 24.32 98.38 98.92 88.65 90.27 10.27 28.11 98.92 100.0 27.57 25.41 02.7 03.24
it 86.49 90.27 65.95 69.19 67.57 72.97 98.38 99.46 92.97 93.51 70.81 90.27 89.19 95.14 71.35 70.27 65.41 69.19
ta 39.46 34.05 00.0 00.0 00.0 00.0 43.78 32.43 15.68 16.76 00.54 01.62 81.62 76.22 00.0 00.0 00.0 00.0
th 99.46 99.46 12.43 18.92 05.95 03.78 81.08 77.84 24.32 27.03 02.16 14.05 100.0 100.0 02.7 02.7 01.08 01.08
tr 74.05 74.05 47.57 48.65 35.14 43.78 92.43 89.19 74.59 077.3 077.3 86.49 72.43 87.03 55.14 56.76 38.92 48.11
LAPE
en 91.35 90.27 100.0 100.0 49.73 21.08 100.0 98.92 100.0 100.0 95.14 22.16 97.84 97.84 097.3 97.84 80.54 07.57
nl 95.68 92.97 082.7 78.38 49.19 38.92 99.46 96.76 100.0 100.0 75.68 65.41 100.0 99.46 88.11 83.78 42.16 30.81
id 95.68 96.22 63.78 61.62 57.84 34.59 99.46 95.14 99.46 99.46 45.95 14.05 99.46 96.22 90.27 91.35 44.86 17.84
ms 92.43 95.68 52.43 44.32 33.51 017.3 99.46 98.92 99.46 99.46 53.51 39.46 98.38 96.76 71.35 71.35 44.32 28.65
vi 100.0 100.0 88.11 080.0 14.59 01.08 100.0 100.0 99.46 99.46 18.92 05.95 100.0 100.0 91.89 88.11 037.3 07.57
jp 100.0 100.0 60.54 49.19 08.11 00.54 100.0 100.0 95.68 95.68 13.51 03.24 100.0 100.0 32.97 32.43 08.11 00.0
zh 100.0 100.0 49.19 042.7 15.14 07.57 100.0 100.0 96.22 94.05 15.68 04.32 100.0 100.0 062.7 55.14 40.54 00.0
fr 98.92 98.92 97.84 92.97 52.97 05.41 99.46 99.46 100.0 100.0 32.97 16.76 99.46 99.46 99.46 98.38 65.95 06.49
pt 98.92 98.92 90.27 88.65 44.86 10.27 100.0 100.0 98.38 98.38 49.73 21.62 99.46 98.92 94.05 91.89 67.03 14.05
ru 100.0 100.0 81.62 077.3 03.24 02.7 100.0 100.0 99.46 99.46 09.19 02.16 100.0 100.0 91.35 81.62 01.08 00.0
it 97.84 98.38 87.57 87.03 64.32 14.59 100.0 97.84 100.0 100.0 55.14 26.49 99.46 99.46 98.38 90.81 71.89 16.22
ta 100.0 100.0 00.0 00.0 00.0 00.0 100.0 99.46 85.41 85.95 00.0 00.0 100.0 100.0 00.0 00.0 00.54 00.0
th 100.0 100.0 52.43 53.51 00.0 00.0 100.0 100.0 85.41 84.86 00.0 00.0 100.0 100.0 41.08 18.38 00.54 00.0
tr 98.92 99.46 64.86 60.54 52.43 44.86 74.59 61.08 96.22 95.68 58.38 55.14 99.46 99.46 71.35 70.27 59.46 51.89
Table 6: Language Steering Shift (LSS) scores from et to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 98.38 97.84 92.43 95.14 43.24 69.73 98.92 99.46 58.92 58.92 49.73 94.59 98.38 98.38 72.97 73.51 54.05 51.89
nl 76.76 78.92 56.22 57.84 35.14 47.03 44.32 32.97 43.24 45.95 57.84 70.81 88.11 96.22 58.38 56.22 25.41 27.03
id 75.14 78.92 61.08 61.62 34.59 49.73 88.65 85.95 81.08 080.0 68.65 84.86 95.68 98.38 57.84 52.97 017.3 15.68
ms 69.19 69.19 057.3 56.22 31.89 39.46 47.57 037.3 38.92 040.0 40.54 58.38 75.14 93.51 042.7 42.16 14.05 14.59
vi 69.73 78.92 51.35 53.51 30.27 40.54 81.08 74.05 47.03 43.78 08.11 14.05 97.84 097.3 45.95 48.11 07.03 11.89
jp 58.38 60.0 18.92 23.24 11.35 14.59 76.76 87.03 38.38 38.92 31.35 48.11 100.0 100.0 15.68 16.76 02.7 01.62
zh 59.46 94.05 24.32 24.86 17.84 15.68 70.27 92.97 26.49 24.86 29.19 042.7 96.76 100.0 14.59 14.05 04.32 01.08
pt 89.19 88.11 46.49 41.08 25.41 38.38 95.68 97.84 75.68 75.68 66.49 77.84 90.81 91.89 45.95 49.19 10.27 20.54
ru 90.81 95.68 32.43 31.35 10.27 13.51 90.81 92.97 28.11 27.57 05.95 08.11 97.84 100.0 14.59 13.51 01.62 02.16
et 100.0 99.46 61.08 58.38 36.76 46.49 50.81 63.24 33.51 30.81 23.78 40.54 100.0 99.46 52.43 51.35 14.05 017.3
it 80.54 72.43 61.62 060.0 51.35 56.22 89.19 96.76 72.43 70.27 75.14 83.24 95.14 98.38 54.05 54.05 36.22 29.73
ta 45.95 45.41 00.54 00.54 00.0 00.0 8.11 02.16 02.16 02.7 00.0 00.0 63.78 60.54 00.0 00.0 00.0 00.0
th 91.35 91.89 18.38 14.05 03.78 06.49 037.3 49.73 08.65 08.11 00.0 03.78 99.46 100.0 06.49 07.57 01.08 00.54
tr 80.0 79.46 58.38 58.92 35.14 42.16 48.11 042.7 23.78 23.78 38.38 42.16 75.68 98.92 36.22 37.84 10.27 10.81
LAPE
en 100.0 100.0 097.3 90.27 92.43 10.27 100.0 100.0 100.0 100.0 097.3 52.97 100.0 100.0 99.46 99.46 100.0 17.84
nl 100.0 100.0 74.59 61.62 060.0 27.03 99.46 96.76 88.65 88.65 64.32 37.84 100.0 98.92 94.59 88.65 68.65 10.81
id 100.0 100.0 70.27 68.11 78.92 11.89 100.0 92.43 98.38 99.46 73.51 58.92 99.46 97.84 78.38 76.76 70.27 09.19
ms 100.0 100.0 55.14 51.89 42.16 12.97 100.0 96.22 98.38 94.59 39.46 23.24 100.0 100.0 70.27 71.35 31.35 04.32
vi 100.0 100.0 55.68 54.05 14.59 05.41 100.0 98.92 88.11 87.03 16.76 08.65 100.0 100.0 93.51 84.86 017.3 08.11
jp 100.0 100.0 47.03 44.32 10.27 00.54 100.0 100.0 56.22 56.76 30.81 29.73 100.0 100.0 18.92 19.46 10.27 00.0
zh 100.0 100.0 26.49 24.32 15.68 00.0 98.92 100.0 59.46 66.49 18.38 03.24 100.0 100.0 34.59 30.81 042.7 02.7
pt 100.0 100.0 93.51 90.81 72.43 12.43 100.0 100.0 96.76 097.3 080.0 40.54 99.46 98.92 91.35 84.86 54.05 16.22
ru 100.0 100.0 79.46 61.08 03.24 00.0 100.0 100.0 89.73 89.19 05.41 00.0 100.0 100.0 54.05 040.0 01.08 00.54
et 100.0 100.0 53.51 48.65 72.43 41.08 100.0 94.59 99.46 98.92 48.65 13.51 100.0 100.0 61.08 60.54 81.08 24.86
it 100.0 100.0 66.49 65.41 74.05 07.57 96.76 95.14 097.3 97.84 76.76 63.24 100.0 100.0 85.95 80.54 057.3 34.05
ta 100.0 100.0 00.54 00.0 00.0 00.0 100.0 89.19 31.89 32.97 00.0 00.0 100.0 100.0 00.0 00.0 00.0 00.0
th 100.0 100.0 43.78 44.32 00.54 00.0 100.0 097.3 60.54 51.89 01.08 00.54 100.0 100.0 16.76 05.41 00.0 00.0
tr 100.0 100.0 52.97 50.81 49.73 14.05 75.14 077.3 96.76 96.22 65.41 042.7 100.0 100.0 51.35 51.35 35.68 05.41
Table 7: Language Steering Shift (LSS) scores from fr to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 96.22 95.14 94.05 93.51 75.68 88.65 98.92 99.46 64.86 66.49 46.49 85.95 98.38 99.46 91.89 92.43 55.14 57.84
nl 90.27 94.05 78.38 80.54 56.22 79.46 40.54 38.38 50.27 48.11 47.57 68.65 94.59 97.3 080.0 77.84 25.95 23.78
id 90.81 89.73 74.59 70.27 48.65 71.89 73.51 71.89 76.22 76.76 76.76 87.03 96.22 98.38 67.03 67.03 22.16 15.14
ms 72.43 74.59 53.51 54.05 38.38 39.46 34.59 29.19 040.0 39.46 38.92 54.05 87.03 92.97 36.22 040.0 14.59 11.89
vi 90.27 93.51 68.11 68.11 44.32 66.49 80.54 73.51 52.97 54.05 12.43 21.62 99.46 98.92 70.81 69.73 07.03 09.19
jp 86.49 91.35 48.11 43.24 12.97 16.76 62.16 78.38 48.11 47.03 32.43 042.7 100.0 98.92 35.68 26.49 03.24 01.62
zh 85.95 95.68 30.27 28.65 29.19 27.57 63.78 90.27 26.49 24.86 25.95 33.51 097.3 100.0 18.92 16.76 09.19 01.08
fr 94.59 96.76 83.24 81.08 79.46 82.16 88.11 92.97 71.89 71.89 51.89 73.51 097.3 97.84 81.62 82.16 29.73 46.49
pt 78.38 75.68 68.11 71.89 51.89 062.7 84.86 82.16 76.76 74.05 63.78 84.32 78.92 83.24 060.0 62.16 022.7 24.32
ru 94.59 96.22 45.41 47.57 34.59 33.51 89.19 90.27 31.89 31.35 08.11 13.51 97.84 99.46 23.78 21.08 01.08 02.16
et 92.97 94.59 71.89 72.97 49.73 58.92 55.14 52.97 36.22 35.68 18.38 35.68 98.38 98.92 56.22 56.22 13.51 23.78
ta 54.59 54.59 01.62 01.62 00.0 00.0 8.11 02.16 01.62 01.08 00.0 00.0 76.22 71.89 00.54 00.54 00.0 00.0
th 98.38 98.38 18.38 20.54 07.57 08.11 34.05 46.49 11.89 14.05 01.62 07.03 100.0 100.0 05.95 05.41 01.08 01.08
tr 85.41 082.7 53.51 56.22 42.16 46.49 29.73 29.19 31.89 27.57 38.38 35.14 93.51 97.3 38.92 43.24 12.97 14.05
LAPE
en 100.0 98.92 94.05 93.51 74.59 22.16 100.0 100.0 100.0 100.0 97.84 06.49 99.46 98.38 95.68 95.14 92.97 14.59
nl 100.0 100.0 87.03 86.49 74.59 037.3 100.0 96.76 98.92 98.92 67.03 022.7 100.0 100.0 97.84 95.68 61.08 11.89
id 100.0 100.0 78.92 70.81 76.76 10.81 99.46 93.51 99.46 99.46 74.05 38.38 100.0 100.0 88.11 94.05 65.41 03.24
ms 100.0 100.0 49.19 48.65 41.08 15.14 99.46 097.3 99.46 99.46 38.92 21.08 100.0 100.0 70.81 69.19 23.24 06.49
vi 100.0 100.0 080.0 78.92 21.08 03.24 100.0 98.92 98.38 93.51 21.08 12.97 100.0 100.0 91.89 88.65 18.38 05.95
jp 100.0 100.0 71.35 66.49 21.62 01.62 100.0 100.0 48.65 53.51 25.95 09.19 100.0 100.0 18.92 21.08 11.89 00.0
zh 100.0 100.0 41.08 39.46 16.22 04.86 98.92 100.0 55.14 58.92 24.86 03.78 100.0 100.0 040.0 36.76 48.65 00.54
fr 100.0 100.0 95.68 94.05 59.46 17.84 100.0 100.0 097.3 96.22 81.62 47.57 100.0 99.46 97.84 96.76 64.86 32.43
pt 98.92 99.46 86.49 83.24 52.43 05.41 100.0 100.0 96.76 96.76 85.95 52.97 100.0 100.0 92.43 90.27 56.22 25.95
ru 100.0 100.0 86.49 78.92 05.41 01.62 100.0 100.0 92.97 93.51 12.43 02.16 100.0 100.0 83.78 55.14 02.7 00.0
et 100.0 98.92 74.05 057.3 74.05 44.32 100.0 95.14 99.46 99.46 61.08 020.0 99.46 99.46 85.41 84.32 63.78 14.05
ta 100.0 100.0 00.54 00.54 00.0 00.0 100.0 93.51 27.03 26.49 00.0 00.0 100.0 100.0 00.54 00.54 00.0 00.0
th 100.0 100.0 44.86 45.95 01.08 00.0 100.0 100.0 61.62 51.35 01.62 01.62 100.0 100.0 12.43 10.27 00.54 00.0
tr 98.92 100.0 49.19 51.89 51.35 15.68 89.73 82.16 98.38 96.22 060.0 35.68 100.0 100.0 49.73 49.19 31.35 10.27
Table 8: Language Steering Shift (LSS) scores from it to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 100.0 100.0 100.0 100.0 95.14 96.22 100.0 98.92 97.84 097.3 58.92 91.35 100.0 100.0 100.0 99.46 94.05 91.35
nl 99.46 98.92 96.22 097.3 60.54 72.43 89.73 92.43 69.19 73.51 52.97 65.41 100.0 100.0 92.43 92.97 14.59 17.84
id 100.0 99.46 92.43 94.05 38.38 67.57 95.68 92.97 95.14 93.51 58.38 79.46 100.0 100.0 90.27 90.27 16.76 16.22
ms 99.46 99.46 75.14 75.68 32.97 040.0 76.76 71.35 63.78 62.16 28.65 52.97 100.0 99.46 75.14 75.68 12.97 09.73
vi 100.0 100.0 097.3 97.84 78.38 83.24 95.68 94.05 76.76 76.76 06.49 21.08 100.0 100.0 87.03 89.73 16.76 18.92
zh 81.62 88.11 75.14 77.84 27.57 53.51 82.16 91.89 59.46 58.38 56.22 68.11 91.35 91.89 057.3 58.92 24.86 32.97
fr 98.92 98.38 92.97 92.43 69.73 77.84 96.22 98.38 86.49 86.49 44.32 78.38 100.0 100.0 91.89 91.89 36.22 40.54
pt 99.46 100.0 97.84 097.3 38.92 64.86 100.0 100.0 84.32 87.03 50.81 80.54 99.46 100.0 93.51 92.43 14.59 10.81
ru 98.92 99.46 69.19 69.19 49.19 54.59 88.11 89.73 56.76 56.76 24.86 38.92 99.46 100.0 78.92 79.46 15.14 14.05
et 100.0 100.0 94.59 96.22 38.92 68.65 75.68 76.22 54.05 52.43 07.03 020.0 100.0 100.0 85.95 87.03 15.68 14.05
it 99.46 96.22 93.51 94.59 56.22 78.38 96.22 100.0 84.32 84.86 58.92 85.95 99.46 96.76 91.35 91.89 36.22 27.57
ta 77.84 077.3 01.08 01.08 00.0 00.0 13.51 02.16 06.49 05.95 00.0 00.54 75.14 67.57 01.08 01.08 00.0 00.0
th 99.46 100.0 32.97 037.3 24.32 32.43 38.38 34.59 14.59 14.59 06.49 11.89 100.0 100.0 44.32 44.32 12.43 12.43
tr 98.92 91.35 080.0 79.46 34.05 44.32 82.7 77.84 45.95 44.86 18.92 34.59 100.0 100.0 70.27 70.27 18.38 38.92
LAPE
en 94.05 92.97 100.0 100.0 74.05 26.49 100.0 100.0 100.0 100.0 79.46 07.03 92.43 90.81 100.0 100.0 75.14 17.84
nl 100.0 99.46 96.22 96.22 27.03 12.97 100.0 96.76 100.0 100.0 50.27 09.19 100.0 100.0 98.38 097.3 69.73 24.86
id 100.0 100.0 96.76 96.22 74.05 09.19 99.46 98.38 99.46 99.46 41.62 08.65 100.0 100.0 98.92 99.46 78.92 18.38
ms 100.0 99.46 75.14 74.59 35.68 12.97 99.46 98.92 100.0 100.0 26.49 07.57 100.0 100.0 75.14 76.76 39.46 11.89
vi 100.0 100.0 98.38 98.38 54.59 43.78 100.0 097.3 99.46 99.46 14.59 05.95 100.0 100.0 097.3 87.03 27.03 15.68
zh 92.97 96.22 78.38 79.46 32.97 30.27 88.65 94.05 87.03 87.03 67.03 51.89 92.97 92.43 76.76 75.68 74.59 52.43
fr 98.92 98.92 97.84 93.51 50.81 18.92 100.0 96.76 100.0 100.0 66.49 20.54 100.0 100.0 98.92 94.59 44.32 21.08
pt 99.46 98.92 98.92 97.84 56.76 17.84 100.0 99.46 100.0 100.0 64.86 21.62 100.0 99.46 90.81 89.73 58.38 07.03
ru 100.0 100.0 91.89 92.97 022.7 20.54 100.0 98.92 097.3 96.76 23.24 08.65 100.0 100.0 93.51 89.19 32.43 28.65
et 100.0 100.0 96.22 90.27 057.3 022.7 98.92 95.14 100.0 100.0 36.76 09.19 100.0 100.0 95.68 95.68 59.46 21.08
it 98.38 98.92 94.59 92.43 78.92 16.76 100.0 98.38 100.0 100.0 61.62 022.7 99.46 98.92 097.3 96.22 54.59 13.51
ta 100.0 100.0 01.62 01.62 00.0 00.0 75.14 12.97 55.68 56.76 01.08 00.54 100.0 99.46 00.54 00.54 00.0 00.0
th 100.0 100.0 61.08 61.08 09.73 07.57 100.0 81.08 060.0 58.92 07.03 05.95 100.0 100.0 042.7 44.32 08.65 05.95
tr 100.0 100.0 73.51 69.73 24.32 05.41 94.59 082.7 99.46 99.46 38.38 20.54 100.0 100.0 74.05 75.14 24.86 17.84
Table 9: Language Steering Shift (LSS) scores from jp to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 100.0 100.0 98.92 98.92 93.51 96.76 96.22 97.84 92.97 93.51 41.08 67.03 100.0 99.46 98.38 98.92 52.43 51.89
nl 96.22 96.22 83.78 83.78 81.62 077.3 82.16 77.84 73.51 74.05 037.3 55.68 097.3 99.46 81.62 84.86 42.16 48.11
id 61.08 46.49 67.57 67.03 70.81 71.35 71.89 74.05 69.73 69.73 57.84 62.16 65.41 49.19 66.49 67.03 64.32 69.19
vi 96.76 98.38 87.57 88.11 72.43 71.35 79.46 76.76 54.05 57.84 16.76 43.24 99.46 99.46 77.84 080.0 18.92 34.59
jp 92.97 94.59 70.81 72.97 35.68 35.68 85.41 88.65 68.65 69.73 29.19 38.92 100.0 100.0 057.3 55.68 06.49 05.95
zh 94.59 99.46 60.54 63.24 53.51 042.7 89.19 92.43 68.11 67.57 27.57 39.46 100.0 100.0 36.76 037.3 05.95 00.0
fr 95.14 94.05 85.41 88.65 82.16 84.32 92.97 93.51 80.54 80.54 45.95 61.62 98.92 99.46 88.65 87.57 54.05 52.43
pt 88.11 86.49 69.73 69.73 68.11 65.41 097.3 97.84 86.49 85.95 48.65 65.95 95.68 98.38 65.95 68.65 32.43 47.03
ru 96.76 98.38 65.95 63.78 41.08 43.24 93.51 94.59 69.73 68.65 04.86 11.35 100.0 100.0 042.7 41.62 03.24 07.03
et 96.22 95.68 83.24 84.32 077.3 78.92 85.41 89.19 68.11 64.86 15.14 45.41 97.84 98.38 83.78 82.16 35.68 037.3
it 97.3 93.51 85.41 87.57 080.0 83.24 95.14 96.76 84.32 84.86 56.76 69.73 94.59 97.3 080.0 77.84 54.05 55.68
ta 82.16 080.0 02.16 02.16 00.0 00.0 36.76 32.43 30.81 28.65 00.0 00.54 84.86 83.78 00.0 00.0 00.0 00.0
th 100.0 100.0 35.14 34.59 18.38 017.3 55.68 59.46 25.95 28.11 00.54 01.08 100.0 100.0 13.51 14.05 01.08 01.62
tr 98.92 98.92 83.24 82.16 53.51 67.03 76.22 83.24 63.24 63.24 43.78 67.57 98.38 98.92 77.84 77.84 020.0 35.14
LAPE
en 100.0 98.92 100.0 100.0 92.97 27.03 100.0 98.92 100.0 99.46 95.68 23.24 100.0 100.0 99.46 100.0 94.05 18.38
nl 100.0 100.0 90.27 88.11 67.03 52.43 100.0 100.0 100.0 100.0 68.11 26.49 100.0 100.0 097.3 96.76 65.95 12.43
id 52.43 50.81 64.86 64.86 60.54 060.0 44.32 38.92 60.54 61.08 042.7 43.78 65.41 69.19 64.86 64.86 54.05 52.97
vi 100.0 100.0 90.81 87.03 29.19 17.84 100.0 99.46 91.89 92.97 19.46 10.27 100.0 100.0 97.84 92.43 30.27 20.54
jp 100.0 100.0 077.3 75.68 23.24 00.0 100.0 100.0 82.16 83.24 10.27 05.95 100.0 100.0 50.81 49.73 11.35 00.54
zh 100.0 100.0 72.43 63.78 12.97 06.49 100.0 100.0 097.3 96.22 37.84 05.95 100.0 100.0 57.84 54.05 42.16 02.7
fr 99.46 99.46 99.46 98.38 81.62 16.22 100.0 100.0 95.68 097.3 65.95 38.92 100.0 100.0 100.0 96.76 87.03 27.03
pt 097.3 97.84 98.92 95.14 72.97 21.08 100.0 100.0 097.3 097.3 082.7 48.65 100.0 100.0 100.0 95.14 83.24 16.76
ru 100.0 100.0 95.14 93.51 02.7 01.62 100.0 100.0 097.3 98.38 16.76 02.7 100.0 100.0 91.89 83.24 05.41 01.08
et 96.22 94.05 86.49 85.95 077.3 67.03 99.46 88.65 100.0 98.92 55.68 34.05 100.0 98.92 88.11 88.11 54.59 34.05
it 98.38 98.38 95.14 94.05 080.0 35.14 100.0 100.0 96.22 94.59 73.51 40.54 99.46 99.46 99.46 96.22 85.95 25.41
ta 100.0 100.0 02.16 02.16 00.0 00.0 100.0 89.19 75.14 77.84 00.0 00.0 100.0 100.0 01.62 01.62 00.0 00.0
th 100.0 100.0 67.03 64.86 00.54 00.54 100.0 100.0 69.19 69.73 01.08 00.0 100.0 100.0 48.65 41.08 00.0 00.0
tr 100.0 100.0 88.65 84.32 46.49 30.81 077.3 077.3 94.59 89.73 54.05 46.49 99.46 99.46 85.95 85.41 35.68 19.46
Table 10: Language Steering Shift (LSS) scores from ms to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 99.46 99.46 99.46 100.0 84.32 96.76 98.38 99.46 81.08 082.7 057.3 082.7 100.0 98.92 99.46 98.92 55.14 75.68
id 94.05 95.68 56.22 57.84 54.59 057.3 94.05 90.81 89.73 90.81 68.65 87.03 93.51 95.68 53.51 43.78 31.89 35.14
ms 080.0 87.57 55.68 057.3 040.0 49.19 68.65 60.54 57.84 56.22 54.59 59.46 062.7 86.49 45.95 042.7 18.38 21.62
vi 88.11 90.27 80.54 78.92 60.54 67.03 91.89 86.49 65.41 70.81 10.27 24.86 100.0 99.46 71.89 71.89 11.35 16.76
jp 94.59 95.14 42.16 39.46 18.92 16.76 89.19 89.73 55.68 057.3 31.35 44.32 98.38 98.38 22.16 21.62 02.7 01.62
zh 89.19 96.76 38.92 037.3 43.24 25.95 90.81 97.3 45.95 45.95 31.89 46.49 100.0 100.0 14.05 15.14 07.03 01.62
fr 86.49 89.19 74.05 75.14 80.54 77.84 95.68 94.59 83.78 83.78 63.24 78.92 94.59 98.92 76.22 78.38 32.43 48.11
pt 90.27 93.51 72.97 74.59 68.65 70.81 097.3 98.38 90.27 90.81 59.46 85.95 91.89 93.51 74.59 74.59 020.0 31.89
ru 95.68 98.38 40.54 040.0 25.41 28.11 96.22 95.68 54.05 52.43 16.22 19.46 99.46 100.0 21.08 23.24 03.78 03.78
et 95.14 95.68 49.19 49.73 48.65 51.89 68.65 64.32 49.19 50.27 29.73 43.24 98.92 98.92 47.03 47.57 32.43 35.14
it 97.84 95.68 83.24 85.95 68.11 74.05 98.38 99.46 83.24 84.86 64.86 84.86 97.3 96.76 77.84 75.68 56.22 50.81
ta 45.95 042.7 00.0 00.0 00.0 00.0 23.78 17.84 13.51 13.51 00.0 00.0 81.62 64.32 00.0 00.0 00.0 00.0
th 99.46 100.0 16.22 16.76 10.81 11.89 52.97 63.24 16.22 16.76 01.62 07.03 100.0 100.0 14.05 14.59 01.08 01.08
tr 85.41 91.35 52.97 52.97 042.7 46.49 76.22 57.84 40.54 042.7 52.43 47.57 92.97 96.76 50.27 52.43 23.78 27.03
LAPE
en 98.92 99.46 98.92 98.92 81.62 017.3 100.0 99.46 99.46 98.92 96.76 03.24 95.14 95.14 97.84 99.46 94.05 07.57
id 98.92 99.46 63.24 56.22 81.08 54.59 100.0 98.92 100.0 100.0 67.57 41.62 99.46 99.46 86.49 91.35 64.32 24.86
ms 100.0 100.0 54.59 52.43 42.16 020.0 100.0 96.76 100.0 99.46 36.22 14.59 99.46 100.0 062.7 62.16 26.49 12.97
vi 100.0 100.0 90.27 87.57 18.38 02.7 100.0 98.38 97.84 98.92 17.84 05.95 100.0 100.0 91.35 85.41 29.19 08.11
jp 100.0 100.0 75.68 72.43 10.27 00.0 100.0 100.0 082.7 86.49 29.73 09.19 100.0 100.0 17.84 20.54 11.89 01.62
zh 100.0 100.0 36.76 36.22 11.89 00.54 99.46 100.0 91.35 90.27 34.59 10.27 100.0 100.0 44.86 36.76 49.19 03.78
fr 98.38 097.3 92.43 90.27 61.62 35.14 100.0 98.92 95.68 95.14 062.7 21.62 100.0 100.0 98.38 96.76 75.68 15.14
pt 100.0 100.0 097.3 95.14 080.0 07.03 100.0 100.0 94.05 94.05 78.38 21.62 100.0 100.0 97.84 94.59 80.54 03.24
ru 100.0 100.0 83.78 080.0 03.24 00.54 100.0 100.0 96.22 96.22 23.24 11.35 100.0 100.0 91.35 69.73 03.78 01.08
et 98.92 98.38 63.78 50.27 52.97 46.49 95.68 81.62 98.38 97.84 34.59 022.7 100.0 100.0 81.08 81.62 63.24 49.19
it 98.92 99.46 85.95 85.41 077.3 26.49 98.92 97.84 98.38 98.92 63.24 21.08 100.0 100.0 94.05 96.22 082.7 12.43
ta 100.0 100.0 01.08 01.08 00.0 00.0 100.0 98.92 52.97 56.76 00.0 00.0 100.0 100.0 01.08 01.08 00.0 00.0
th 100.0 100.0 33.51 34.59 00.54 00.0 100.0 100.0 83.24 83.24 01.08 00.0 100.0 100.0 25.95 21.08 00.0 00.0
tr 100.0 100.0 60.54 61.08 44.86 23.78 91.89 81.62 90.81 89.73 060.0 040.0 100.0 100.0 57.84 57.84 44.32 23.24
Table 11: Language Steering Shift (LSS) scores from nl to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 99.46 99.46 100.0 100.0 72.97 95.14 96.76 96.22 68.65 65.95 52.43 84.86 100.0 100.0 99.46 98.92 65.41 65.41
nl 97.84 097.3 87.57 88.11 63.24 69.73 54.05 040.0 48.11 48.65 61.62 70.27 98.92 100.0 80.54 080.0 47.03 47.03
id 97.84 99.46 71.35 68.65 56.22 59.46 76.22 58.38 72.97 72.97 71.35 88.65 100.0 100.0 72.43 70.81 38.38 29.19
ms 81.08 96.22 56.76 56.76 45.95 47.03 36.22 26.49 38.92 37.84 37.84 55.68 91.89 98.92 47.57 47.03 23.24 020.0
vi 88.65 93.51 58.38 58.38 51.89 53.51 81.08 75.68 49.19 50.27 08.11 16.22 100.0 100.0 077.3 78.38 022.7 022.7
jp 94.59 95.14 44.32 47.03 16.76 24.32 69.19 82.16 45.41 44.32 30.27 040.0 100.0 100.0 25.95 27.03 03.24 02.16
zh 90.27 97.3 040.0 50.27 28.11 24.32 56.22 82.7 21.08 017.3 23.78 040.0 99.46 100.0 22.16 23.24 12.97 00.0
fr 93.51 96.22 79.46 080.0 83.78 83.24 88.11 90.81 74.59 74.05 58.92 76.76 98.92 100.0 80.54 81.62 53.51 59.46
ru 98.38 100.0 52.97 58.38 23.78 30.81 78.92 84.32 23.78 21.62 10.81 10.27 100.0 100.0 25.95 26.49 01.62 03.24
et 97.84 97.84 66.49 71.35 52.43 58.92 54.05 60.54 41.62 35.68 25.95 29.19 99.46 99.46 65.41 64.86 30.27 36.22
it 90.81 88.65 78.38 78.92 67.03 67.57 88.65 87.03 65.95 67.03 63.24 70.81 94.59 98.92 082.7 84.32 68.65 64.32
ta 48.11 48.11 01.62 01.62 00.0 00.0 8.65 04.86 03.24 03.24 00.0 00.0 74.59 66.49 00.0 00.0 00.0 00.0
th 98.38 98.38 017.3 18.38 11.35 12.97 040.0 56.76 04.86 07.57 02.7 03.24 100.0 100.0 12.97 12.97 02.16 01.08
tr 89.19 90.81 62.16 61.08 48.11 52.43 35.14 28.65 23.24 022.7 39.46 36.22 99.46 100.0 46.49 48.11 29.73 31.89
LAPE
en 100.0 98.38 100.0 100.0 082.7 30.27 100.0 100.0 100.0 100.0 99.46 28.65 097.3 95.68 98.92 100.0 91.89 27.03
nl 100.0 100.0 97.84 96.22 71.35 08.11 100.0 98.92 91.89 89.19 68.65 41.62 100.0 100.0 98.38 097.3 74.59 07.03
id 100.0 100.0 94.05 92.43 73.51 29.73 100.0 96.76 95.68 96.22 71.89 45.95 100.0 100.0 99.46 99.46 63.78 08.65
ms 100.0 100.0 51.35 54.05 47.03 18.92 96.76 95.14 92.97 93.51 45.41 30.81 100.0 100.0 72.43 73.51 25.95 01.62
vi 100.0 100.0 73.51 61.62 21.62 08.65 100.0 99.46 92.43 89.73 19.46 12.97 100.0 100.0 91.35 86.49 24.86 10.27
jp 100.0 100.0 78.38 69.19 017.3 01.08 100.0 100.0 54.59 54.59 36.76 10.27 100.0 100.0 29.73 29.19 14.05 00.54
zh 100.0 100.0 41.62 037.3 22.16 02.16 99.46 99.46 63.78 65.41 35.14 07.03 100.0 100.0 39.46 31.35 51.35 05.95
fr 100.0 100.0 96.76 94.05 77.84 29.73 99.46 99.46 95.68 96.22 77.84 59.46 100.0 100.0 99.46 99.46 70.81 30.81
ru 100.0 100.0 85.95 077.3 03.24 03.24 100.0 100.0 90.81 91.89 11.35 01.08 100.0 100.0 83.24 63.24 02.16 02.16
et 99.46 99.46 75.68 55.68 73.51 24.32 100.0 93.51 99.46 98.38 71.35 51.89 100.0 100.0 87.57 88.11 65.41 13.51
it 97.3 97.3 89.19 85.41 75.14 26.49 100.0 99.46 92.43 91.89 82.16 74.05 100.0 100.0 97.84 94.59 73.51 47.57
ta 100.0 100.0 00.54 00.54 00.0 00.0 100.0 097.3 35.14 35.68 00.0 00.0 100.0 100.0 01.08 01.08 00.0 00.0
th 100.0 100.0 52.97 57.84 01.08 00.0 100.0 100.0 66.49 51.89 02.16 02.16 100.0 100.0 24.32 12.97 01.08 00.0
tr 100.0 100.0 62.16 62.16 57.84 32.97 080.0 85.41 96.76 95.14 74.05 54.05 100.0 100.0 71.35 69.19 35.68 04.86
Table 12: Language Steering Shift (LSS) scores from pt to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 100.0 100.0 100.0 100.0 080.0 94.05 100.0 100.0 97.84 97.84 53.51 75.14 100.0 100.0 99.46 100.0 70.81 54.59
nl 99.46 100.0 95.14 94.59 30.81 65.95 96.22 94.59 84.86 86.49 49.19 67.03 100.0 100.0 93.51 92.97 19.46 19.46
id 100.0 100.0 96.76 96.76 30.81 55.68 96.76 94.05 94.59 93.51 64.86 87.57 100.0 100.0 90.81 90.81 21.62 15.68
ms 100.0 100.0 78.38 077.3 24.86 47.57 75.68 68.65 68.11 68.11 39.46 66.49 100.0 100.0 71.89 71.35 12.97 09.73
vi 99.46 98.38 89.73 91.89 49.73 90.27 91.35 88.65 68.65 69.73 15.68 67.03 100.0 100.0 87.57 86.49 25.95 25.95
jp 95.68 96.22 83.78 83.24 06.49 64.32 91.89 90.81 69.19 69.19 49.73 78.92 100.0 100.0 62.16 057.3 19.46 29.19
zh 96.76 98.38 82.16 86.49 18.38 43.24 92.97 93.51 72.97 74.05 46.49 077.3 100.0 100.0 43.24 040.0 31.35 01.62
fr 100.0 99.46 97.84 97.84 54.05 70.27 100.0 100.0 94.05 94.05 55.14 77.84 100.0 100.0 96.22 96.76 48.11 40.54
pt 98.92 097.3 98.92 98.92 27.03 63.78 100.0 100.0 97.84 98.38 67.57 84.32 98.92 99.46 95.68 94.59 017.3 16.76
et 100.0 100.0 95.14 96.76 23.24 52.43 89.73 62.16 75.68 74.05 19.46 43.24 100.0 99.46 71.35 81.08 14.05 16.76
it 98.92 98.92 98.38 98.38 31.89 69.73 99.46 100.0 95.14 96.22 61.08 76.76 100.0 100.0 96.76 96.76 51.35 42.16
ta 87.03 84.32 02.7 03.24 00.54 01.08 12.97 04.86 10.81 11.89 00.54 01.62 94.59 81.08 01.62 01.62 00.0 00.0
th 100.0 100.0 28.65 31.89 29.73 28.11 38.38 38.92 020.0 21.62 11.89 16.76 100.0 100.0 30.81 31.89 09.73 08.65
tr 96.22 91.89 080.0 080.0 27.57 51.35 87.57 76.76 64.32 63.78 45.41 71.89 100.0 100.0 70.81 71.89 14.59 12.43
LAPE
en 100.0 100.0 99.46 99.46 81.08 29.19 100.0 100.0 100.0 100.0 95.14 15.14 100.0 100.0 100.0 100.0 92.43 08.65
nl 100.0 100.0 097.3 96.76 65.41 01.08 99.46 91.89 100.0 100.0 37.84 07.57 100.0 100.0 100.0 98.38 68.65 09.19
id 100.0 100.0 99.46 99.46 84.32 20.54 98.92 92.43 097.3 98.38 62.16 15.14 100.0 100.0 100.0 100.0 73.51 08.65
ms 100.0 100.0 84.32 78.38 46.49 06.49 98.92 97.84 100.0 100.0 47.57 14.59 100.0 100.0 100.0 100.0 34.59 07.57
vi 100.0 100.0 100.0 100.0 29.73 14.59 100.0 98.92 100.0 99.46 19.46 04.86 100.0 100.0 96.22 98.38 28.11 09.19
jp 99.46 99.46 95.14 94.05 16.76 10.27 99.46 097.3 93.51 95.14 28.11 16.76 100.0 100.0 94.59 90.27 11.89 04.32
zh 100.0 100.0 99.46 98.92 28.65 17.84 98.38 98.38 097.3 96.22 48.11 022.7 100.0 100.0 96.76 89.73 50.27 06.49
fr 100.0 100.0 100.0 97.84 71.35 09.73 98.38 98.38 98.38 98.38 69.73 017.3 100.0 100.0 097.3 96.76 69.19 07.03
pt 100.0 100.0 96.22 88.11 73.51 08.11 98.92 99.46 99.46 99.46 65.41 15.68 100.0 100.0 100.0 98.92 75.68 04.32
et 100.0 98.92 97.84 97.84 52.97 34.59 87.03 75.68 99.46 99.46 48.65 19.46 94.59 94.59 96.76 96.76 41.62 28.11
it 100.0 100.0 99.46 98.92 74.59 06.49 89.19 86.49 88.11 86.49 69.19 36.22 100.0 99.46 100.0 100.0 70.81 07.57
ta 98.92 98.38 00.0 01.08 00.54 00.0 62.7 020.0 31.89 25.95 00.54 00.0 99.46 99.46 18.38 15.68 00.0 00.0
th 100.0 100.0 99.46 99.46 02.16 00.0 100.0 79.46 95.68 84.32 03.78 03.78 100.0 100.0 94.05 92.97 00.0 00.0
tr 100.0 100.0 85.95 85.95 30.81 07.03 54.59 56.22 68.11 65.95 31.35 23.78 97.84 97.84 097.3 94.59 35.14 02.16
Table 13: Language Steering Shift (LSS) scores from ru to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.84 100.0 100.0 100.0 100.0 100.0 097.3 96.76
nl 100.0 100.0 100.0 100.0 95.68 100.0 100.0 100.0 98.92 98.92 88.65 100.0 100.0 100.0 100.0 100.0 84.86 81.08
id 100.0 100.0 100.0 100.0 93.51 99.46 100.0 100.0 99.46 98.92 94.59 99.46 100.0 98.38 100.0 100.0 080.0 84.86
ms 100.0 100.0 100.0 100.0 75.68 100.0 100.0 100.0 100.0 99.46 080.0 100.0 100.0 100.0 100.0 100.0 060.0 60.54
vi 100.0 98.92 100.0 100.0 100.0 100.0 99.46 98.92 98.92 98.92 77.84 98.92 100.0 100.0 100.0 100.0 73.51 78.38
jp 97.84 98.92 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.84 100.0 100.0 100.0 100.0 100.0 65.41 96.22
zh 100.0 100.0 100.0 100.0 98.92 100.0 100.0 100.0 98.92 98.38 99.46 100.0 100.0 100.0 100.0 100.0 65.95 25.95
fr 100.0 100.0 100.0 100.0 94.59 100.0 100.0 100.0 100.0 100.0 91.35 100.0 100.0 100.0 100.0 100.0 90.81 83.78
pt 100.0 100.0 100.0 100.0 87.57 100.0 100.0 100.0 100.0 100.0 95.68 100.0 100.0 100.0 100.0 100.0 84.32 81.62
ru 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.92 98.38 88.65 99.46 100.0 98.92 100.0 100.0 23.24 32.43
et 100.0 99.46 100.0 100.0 96.76 100.0 100.0 100.0 98.38 98.92 69.73 99.46 100.0 100.0 100.0 100.0 87.57 92.43
it 100.0 100.0 100.0 100.0 96.76 100.0 100.0 100.0 100.0 100.0 98.38 100.0 100.0 100.0 100.0 100.0 92.43 89.73
th 100.0 100.0 98.38 98.38 98.92 98.92 97.84 98.38 76.22 76.76 87.57 93.51 99.46 100.0 97.84 097.3 24.32 79.46
tr 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.46 98.92 082.7 98.92 100.0 99.46 100.0 100.0 58.92 060.0
LAPE
en 97.84 097.3 100.0 100.0 56.76 12.97 100.0 100.0 100.0 100.0 98.92 21.08 96.22 96.22 100.0 100.0 87.57 24.32
nl 100.0 100.0 100.0 100.0 69.73 32.97 100.0 94.59 100.0 100.0 37.84 07.57 100.0 97.84 100.0 100.0 54.05 11.89
id 100.0 100.0 100.0 100.0 76.76 48.65 100.0 99.46 100.0 100.0 74.05 55.14 100.0 100.0 100.0 100.0 68.11 34.05
ms 100.0 100.0 100.0 100.0 45.95 30.81 100.0 99.46 100.0 100.0 30.81 17.84 99.46 99.46 100.0 100.0 39.46 26.49
vi 100.0 99.46 100.0 100.0 28.11 27.57 99.46 99.46 100.0 100.0 21.62 10.27 100.0 100.0 100.0 100.0 040.0 23.24
jp 100.0 100.0 100.0 100.0 040.0 44.86 100.0 100.0 100.0 100.0 29.73 23.24 100.0 100.0 100.0 100.0 07.57 02.16
zh 100.0 100.0 100.0 100.0 14.05 08.11 100.0 100.0 100.0 100.0 062.7 42.16 100.0 100.0 100.0 100.0 25.41 00.0
fr 99.46 99.46 100.0 100.0 71.35 41.62 100.0 100.0 100.0 100.0 47.03 46.49 100.0 100.0 100.0 100.0 71.89 16.22
pt 100.0 100.0 100.0 100.0 057.3 35.14 100.0 100.0 100.0 100.0 71.35 46.49 100.0 100.0 100.0 100.0 69.73 18.38
ru 100.0 100.0 100.0 100.0 05.41 05.41 100.0 100.0 100.0 100.0 46.49 26.49 100.0 100.0 100.0 100.0 02.16 02.16
et 99.46 99.46 100.0 100.0 88.11 84.86 100.0 097.3 100.0 100.0 057.3 28.65 100.0 98.92 100.0 100.0 54.05 48.65
it 99.46 99.46 100.0 100.0 42.16 11.89 100.0 100.0 100.0 100.0 88.65 88.11 99.46 99.46 100.0 100.0 78.38 21.08
th 100.0 100.0 99.46 99.46 12.97 21.08 100.0 100.0 100.0 99.46 91.35 94.05 100.0 100.0 99.46 99.46 01.62 02.16
tr 100.0 100.0 100.0 100.0 47.03 39.46 100.0 95.68 100.0 100.0 87.57 84.86 100.0 100.0 100.0 100.0 23.78 16.76
Table 14: Language Steering Shift (LSS) scores from ta to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 100.0 100.0 100.0 100.0 98.92 99.46 100.0 100.0 99.46 100.0 97.84 100.0 100.0 100.0 100.0 100.0 65.95 74.05
nl 100.0 100.0 99.46 99.46 88.11 92.43 100.0 97.84 95.68 94.59 81.08 91.89 100.0 100.0 98.92 98.92 32.97 31.89
id 100.0 100.0 99.46 99.46 75.68 94.59 99.46 95.14 99.46 99.46 95.68 99.46 100.0 100.0 99.46 98.38 27.03 30.81
ms 100.0 100.0 97.84 97.84 52.43 70.81 99.46 78.92 94.05 96.22 93.51 99.46 100.0 99.46 97.84 95.68 22.16 33.51
vi 99.46 96.76 99.46 99.46 87.03 99.46 100.0 99.46 98.38 98.38 87.03 96.22 100.0 100.0 99.46 100.0 51.35 69.19
jp 98.92 98.38 91.35 91.89 75.68 87.57 097.3 98.38 94.05 93.51 89.19 95.14 100.0 100.0 93.51 93.51 23.24 50.27
zh 100.0 100.0 97.84 98.38 65.95 64.32 100.0 100.0 97.84 097.3 95.68 98.92 100.0 100.0 87.03 87.03 09.19 05.95
fr 100.0 097.3 100.0 100.0 83.24 89.19 100.0 100.0 99.46 100.0 93.51 100.0 100.0 100.0 100.0 100.0 45.41 38.38
pt 100.0 100.0 100.0 100.0 077.3 81.62 100.0 100.0 100.0 100.0 96.76 100.0 100.0 100.0 99.46 99.46 41.08 32.43
ru 100.0 100.0 95.14 95.14 69.73 88.11 98.92 100.0 92.43 91.35 62.16 92.43 100.0 100.0 95.68 95.68 09.73 14.59
et 100.0 100.0 100.0 100.0 86.49 90.27 99.46 98.92 98.92 98.92 64.86 96.22 100.0 100.0 99.46 99.46 18.92 18.92
it 100.0 100.0 100.0 100.0 85.41 92.97 100.0 100.0 100.0 100.0 91.35 100.0 100.0 100.0 95.68 96.76 45.95 34.59
ta 97.84 93.51 07.03 07.57 04.86 03.24 72.97 52.43 63.78 61.08 56.76 55.14 99.46 98.92 05.95 05.95 00.0 01.08
tr 100.0 100.0 100.0 100.0 72.97 86.49 100.0 95.14 98.38 98.38 87.03 100.0 100.0 100.0 99.46 99.46 24.86 35.14
LAPE
en 100.0 96.22 100.0 100.0 46.49 45.95 100.0 100.0 100.0 100.0 97.84 11.35 100.0 100.0 100.0 100.0 86.49 09.19
nl 100.0 100.0 99.46 99.46 61.62 10.27 100.0 97.84 100.0 100.0 71.89 21.62 100.0 100.0 100.0 100.0 64.32 04.86
id 100.0 100.0 97.84 97.84 11.89 13.51 100.0 99.46 100.0 100.0 61.08 15.68 100.0 100.0 100.0 100.0 58.92 14.59
ms 100.0 100.0 95.14 97.84 12.97 10.27 100.0 100.0 100.0 100.0 28.11 09.19 100.0 100.0 100.0 100.0 25.95 02.16
vi 100.0 100.0 100.0 100.0 08.65 00.0 100.0 100.0 100.0 100.0 020.0 00.54 100.0 100.0 100.0 100.0 19.46 01.08
jp 100.0 100.0 100.0 100.0 10.27 04.86 100.0 100.0 100.0 100.0 23.78 20.54 100.0 100.0 100.0 100.0 04.32 01.08
zh 97.84 85.95 100.0 100.0 20.54 14.59 100.0 100.0 100.0 100.0 51.35 18.92 100.0 100.0 99.46 98.92 61.08 05.95
fr 99.46 99.46 100.0 100.0 51.35 10.27 94.59 94.05 98.92 98.92 73.51 35.14 100.0 100.0 97.84 95.68 077.3 05.41
pt 100.0 100.0 98.92 98.92 78.92 07.03 100.0 100.0 99.46 99.46 62.16 20.54 100.0 100.0 100.0 100.0 83.24 01.62
ru 100.0 100.0 100.0 100.0 37.84 03.78 100.0 100.0 100.0 100.0 060.0 02.16 100.0 100.0 100.0 100.0 11.35 04.32
et 097.3 097.3 100.0 100.0 72.43 47.03 077.3 75.14 90.27 89.73 50.81 34.05 100.0 100.0 99.46 99.46 81.08 042.7
it 100.0 100.0 100.0 100.0 73.51 10.81 96.76 96.22 96.22 95.68 67.03 23.24 100.0 100.0 100.0 100.0 73.51 02.7
ta 99.46 98.92 04.86 04.86 00.0 00.0 91.35 53.51 92.97 90.27 00.54 00.0 100.0 100.0 060.0 56.76 00.0 00.0
tr 100.0 99.46 100.0 100.0 61.62 35.68 76.22 70.81 91.89 91.35 29.73 22.16 100.0 100.0 100.0 100.0 44.86 04.32
Table 15: Language Steering Shift (LSS) scores from th to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 100.0 100.0 100.0 100.0 91.35 100.0 100.0 100.0 97.84 097.3 29.19 77.84 100.0 100.0 100.0 100.0 79.46 64.86
nl 99.46 98.92 84.86 87.03 69.73 077.3 89.73 81.62 83.78 86.49 24.86 077.3 100.0 99.46 84.86 81.62 060.0 53.51
id 95.68 95.68 83.24 85.95 77.84 79.46 95.14 94.05 90.81 89.19 49.73 87.03 97.3 97.3 76.76 78.92 38.92 34.59
ms 84.32 84.86 75.14 74.05 44.32 54.05 66.49 63.24 69.19 69.19 37.84 70.27 89.73 91.35 65.41 64.86 23.78 28.11
vi 96.76 97.3 86.49 87.03 63.24 67.03 97.3 91.35 82.16 082.7 08.65 47.03 99.46 100.0 79.46 080.0 23.78 28.11
jp 94.59 91.35 58.92 62.16 34.05 36.22 90.81 90.27 78.92 080.0 16.22 037.3 98.92 98.92 52.97 50.27 03.24 02.16
zh 90.27 98.38 48.65 54.59 44.86 38.92 97.84 98.92 80.54 080.0 25.41 41.08 100.0 100.0 41.62 38.38 23.78 03.78
fr 98.92 98.38 92.97 92.43 81.62 84.86 99.46 99.46 90.81 90.81 21.62 85.95 98.92 100.0 91.35 91.89 58.38 56.76
pt 97.84 97.84 89.73 86.49 67.03 077.3 100.0 100.0 95.14 95.14 38.92 87.57 94.05 96.76 90.27 91.89 27.03 34.05
ru 98.38 98.38 64.86 63.78 38.38 38.92 097.3 98.92 78.38 73.51 02.7 18.92 99.46 100.0 55.68 55.14 01.62 04.32
et 97.3 94.05 83.24 83.24 78.92 81.62 88.11 71.89 72.43 68.11 08.65 50.27 99.46 98.38 79.46 82.16 61.08 55.14
it 96.76 95.68 94.05 92.97 78.38 83.78 100.0 100.0 93.51 94.05 39.46 91.35 99.46 100.0 93.51 91.35 58.92 55.68
ta 62.7 057.3 00.0 00.0 00.0 00.0 22.7 11.35 11.35 10.81 00.0 00.0 86.49 83.24 00.0 00.0 00.0 00.0
th 100.0 100.0 38.38 41.62 18.92 11.89 81.08 79.46 27.57 27.57 03.78 07.03 100.0 100.0 13.51 10.81 00.54 01.08
LAPE
en 97.84 95.68 100.0 100.0 61.62 34.05 99.46 99.46 100.0 100.0 99.46 23.24 98.92 97.84 100.0 100.0 88.11 16.22
nl 100.0 100.0 93.51 90.81 72.43 25.95 100.0 98.38 100.0 100.0 45.95 06.49 100.0 100.0 100.0 98.38 70.81 38.38
id 98.38 98.92 88.11 84.32 062.7 44.32 99.46 95.14 99.46 99.46 077.3 63.78 100.0 100.0 96.76 097.3 25.41 16.22
ms 98.38 98.92 72.97 75.14 37.84 27.03 100.0 91.35 99.46 98.92 27.57 017.3 98.92 100.0 93.51 92.97 18.92 12.43
vi 100.0 100.0 93.51 89.73 09.73 04.86 100.0 99.46 99.46 99.46 18.92 10.81 100.0 100.0 99.46 96.76 13.51 04.86
jp 100.0 100.0 81.62 76.76 19.46 02.16 100.0 100.0 89.19 89.19 36.22 21.08 100.0 100.0 60.54 060.0 12.97 05.95
zh 100.0 100.0 61.62 53.51 14.05 07.57 100.0 100.0 98.38 99.46 36.76 08.11 100.0 100.0 75.68 65.95 48.11 04.86
fr 98.92 99.46 95.68 95.14 69.73 33.51 100.0 100.0 100.0 99.46 63.78 50.27 100.0 98.92 097.3 097.3 68.11 08.65
pt 97.84 96.76 95.68 95.14 70.27 25.41 100.0 100.0 100.0 100.0 86.49 58.38 100.0 100.0 96.76 92.97 72.97 04.86
ru 100.0 100.0 94.05 88.65 02.7 01.08 100.0 100.0 100.0 100.0 11.89 00.54 100.0 100.0 92.97 76.22 05.41 00.0
et 97.84 097.3 85.41 81.62 85.95 87.03 097.3 83.24 100.0 100.0 037.3 24.32 97.84 96.22 89.19 89.73 082.7 74.05
it 97.84 097.3 96.76 097.3 057.3 11.89 100.0 97.84 99.46 99.46 68.65 36.22 98.38 97.84 98.92 097.3 68.11 09.73
ta 100.0 100.0 01.08 01.08 00.0 00.0 100.0 96.22 80.54 86.49 00.0 00.0 100.0 100.0 00.0 01.08 00.0 00.0
th 100.0 100.0 74.59 71.89 00.0 00.0 100.0 100.0 93.51 94.05 07.03 05.41 100.0 100.0 51.89 29.73 00.54 00.0
Table 16: Language Steering Shift (LSS) scores from tr to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 99.46 100.0 99.46 99.46 93.51 95.68 100.0 100.0 94.59 94.59 69.19 95.68 100.0 100.0 99.46 99.46 77.84 62.16
nl 100.0 100.0 95.14 96.76 64.32 63.78 88.11 78.92 85.95 84.32 59.46 76.22 100.0 100.0 75.14 79.46 40.54 55.14
id 98.38 100.0 67.57 71.89 58.38 61.62 90.27 87.03 89.73 90.81 78.92 94.05 100.0 100.0 76.76 78.92 51.35 59.46
ms 87.57 94.59 63.24 61.62 037.3 38.38 63.78 56.22 56.76 57.84 53.51 72.97 98.92 100.0 48.65 50.27 28.65 38.38
jp 89.19 89.73 52.43 54.05 30.81 17.84 99.46 98.92 78.92 78.92 44.32 81.62 99.46 100.0 40.54 34.59 17.84 15.14
zh 84.32 94.59 50.81 49.73 38.38 04.32 90.81 98.38 83.24 82.16 51.35 75.14 100.0 100.0 017.3 14.05 022.7 08.11
fr 90.81 89.73 83.78 082.7 70.27 74.59 98.38 98.92 90.81 90.27 70.27 90.81 100.0 99.46 83.24 83.24 42.16 52.97
pt 83.78 080.0 82.16 080.0 60.54 65.41 96.76 97.3 93.51 94.59 72.97 91.35 96.22 96.76 082.7 81.08 27.57 36.22
ru 97.84 99.46 63.24 64.32 44.32 44.86 92.97 95.14 70.81 63.78 28.65 47.57 99.46 100.0 51.35 49.73 04.32 06.49
et 98.92 98.92 68.11 67.57 47.57 58.38 88.11 81.62 080.0 81.08 28.65 54.59 100.0 100.0 077.3 77.84 28.11 33.51
it 94.05 90.81 89.19 89.73 44.86 64.86 98.92 99.46 88.65 87.57 75.68 83.24 95.14 96.22 81.62 83.24 39.46 48.65
ta 71.35 71.35 00.0 00.0 00.0 00.0 16.76 03.24 03.24 02.7 00.0 01.08 76.22 76.22 00.0 00.0 00.0 00.0
th 98.38 97.84 017.3 17.84 11.89 08.65 52.97 66.49 23.24 24.86 05.41 15.68 100.0 100.0 18.92 020.0 02.16 02.16
tr 100.0 097.3 65.95 64.32 45.41 47.03 82.7 70.27 58.92 55.14 61.08 80.54 100.0 100.0 47.57 48.65 36.76 47.03
LAPE
en 100.0 100.0 98.38 98.92 92.43 16.22 100.0 100.0 100.0 100.0 99.46 19.46 100.0 100.0 100.0 100.0 097.3 15.14
nl 100.0 100.0 97.84 96.76 79.46 08.11 100.0 100.0 100.0 100.0 67.57 11.35 100.0 100.0 98.38 98.38 84.32 07.03
id 100.0 100.0 89.19 080.0 92.43 88.11 100.0 99.46 99.46 100.0 48.11 35.68 100.0 100.0 98.92 98.38 74.05 47.03
ms 99.46 99.46 52.43 49.19 86.49 64.32 100.0 99.46 100.0 100.0 31.89 21.08 100.0 100.0 90.81 84.86 33.51 022.7
jp 100.0 100.0 062.7 61.62 09.19 05.95 100.0 100.0 98.38 98.92 24.86 10.27 100.0 100.0 48.11 47.03 15.68 09.19
zh 100.0 100.0 39.46 33.51 020.0 01.62 100.0 100.0 93.51 94.59 52.43 14.05 100.0 100.0 61.62 48.65 43.78 03.78
fr 99.46 98.92 98.92 93.51 70.81 21.08 100.0 100.0 99.46 100.0 55.14 16.76 100.0 100.0 100.0 100.0 60.54 11.35
pt 100.0 100.0 94.59 91.89 54.05 16.76 98.92 98.38 99.46 99.46 78.38 18.38 99.46 99.46 96.22 93.51 71.89 03.78
ru 100.0 100.0 85.41 71.35 06.49 03.24 100.0 100.0 100.0 100.0 19.46 01.62 100.0 100.0 87.03 79.46 02.7 00.0
et 100.0 100.0 89.19 84.86 80.54 25.95 99.46 82.16 100.0 100.0 48.65 10.27 100.0 100.0 90.27 90.81 73.51 11.89
it 100.0 100.0 98.38 097.3 63.78 07.57 100.0 100.0 100.0 100.0 87.57 14.05 100.0 100.0 100.0 99.46 83.24 14.59
ta 100.0 100.0 00.0 00.0 00.0 00.0 100.0 90.81 48.65 50.81 00.0 00.0 100.0 100.0 01.62 01.62 00.0 00.0
th 100.0 100.0 32.43 35.14 04.86 03.24 100.0 100.0 79.46 75.68 03.78 02.7 100.0 100.0 24.32 16.22 01.62 01.08
tr 100.0 100.0 61.62 61.08 83.24 71.89 81.08 74.59 100.0 100.0 060.0 52.43 100.0 100.0 67.57 68.11 41.62 32.97
Table 17: Language Steering Shift (LSS) scores from vi to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Qwen Gemma SeaLLM
pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$ pmax $\uparrow$ pmed $\uparrow$ =max $\uparrow$ +max $\uparrow$ =0 $\downarrow$ =10p $\downarrow$
Baseline
en 100.0 100.0 100.0 100.0 84.86 96.76 100.0 100.0 95.14 95.14 51.35 71.35 100.0 100.0 100.0 100.0 58.92 43.24
nl 100.0 100.0 96.76 95.14 38.92 67.57 90.27 92.43 81.08 79.46 35.68 42.16 100.0 100.0 93.51 95.14 05.95 05.41
id 100.0 100.0 95.68 93.51 31.35 56.22 97.3 96.22 95.68 95.68 48.11 75.14 100.0 100.0 91.35 91.35 11.89 05.95
ms 100.0 99.46 77.84 76.22 24.86 45.41 76.76 71.89 70.27 69.73 27.03 54.59 100.0 100.0 69.19 67.57 05.95 04.86
vi 100.0 100.0 98.38 97.84 042.7 85.41 91.89 082.7 67.57 70.81 10.81 17.84 100.0 100.0 88.11 94.05 09.19 06.49
jp 89.19 90.27 72.43 73.51 47.57 58.38 79.46 79.46 63.78 61.08 35.14 63.24 91.35 92.43 65.95 67.03 25.95 26.49
fr 100.0 100.0 93.51 93.51 65.95 73.51 097.3 100.0 90.27 90.27 41.62 59.46 100.0 100.0 94.05 93.51 16.22 09.73
pt 100.0 100.0 98.92 99.46 33.51 65.41 98.92 100.0 92.97 92.97 52.97 67.57 100.0 100.0 96.22 96.22 05.41 03.78
ru 99.46 99.46 76.22 77.84 35.68 46.49 85.41 90.27 062.7 060.0 23.78 32.43 99.46 100.0 80.54 82.16 03.24 03.24
et 100.0 100.0 99.46 98.92 37.84 52.97 89.19 75.14 71.89 67.03 09.19 21.62 100.0 100.0 88.11 91.35 02.7 01.08
it 100.0 99.46 96.76 95.68 51.35 74.05 97.84 100.0 91.35 90.27 36.76 79.46 100.0 100.0 89.73 89.73 21.62 09.19
ta 85.95 85.41 03.24 03.24 00.0 00.0 7.57 03.24 03.78 03.78 00.0 01.62 82.16 69.19 00.0 00.0 00.0 00.0
th 100.0 98.92 042.7 43.78 14.05 19.46 52.43 52.43 14.05 14.59 02.7 07.57 100.0 100.0 32.43 34.05 01.62 01.08
tr 100.0 100.0 78.38 78.92 23.24 48.11 89.19 82.16 64.32 64.86 27.03 35.14 100.0 100.0 70.81 69.19 04.32 03.24
LAPE
en 63.24 72.43 100.0 100.0 27.03 29.73 100.0 99.46 100.0 100.0 90.81 56.76 61.62 65.41 90.81 93.51 45.95 35.68
nl 100.0 100.0 92.97 94.05 40.54 09.73 99.46 89.73 100.0 100.0 46.49 11.35 100.0 100.0 97.84 097.3 057.3 08.65
id 96.76 97.3 87.57 92.97 83.78 32.43 100.0 92.43 99.46 99.46 42.16 09.19 100.0 100.0 95.68 94.05 67.57 07.57
ms 95.68 96.22 65.41 65.95 47.57 23.24 100.0 92.43 100.0 100.0 28.11 07.03 98.38 99.46 70.27 69.19 30.81 03.24
vi 96.22 92.97 97.3 95.68 27.57 18.38 97.3 89.73 94.59 93.51 15.14 07.57 98.38 95.68 96.76 93.51 14.59 05.95
jp 91.35 90.27 76.76 76.22 30.27 31.35 98.92 92.97 84.32 87.03 45.41 41.08 91.89 90.81 60.54 50.27 13.51 10.27
fr 92.43 91.89 95.68 91.35 60.54 27.03 98.92 98.92 100.0 100.0 65.95 41.08 98.38 98.92 94.59 92.97 50.27 08.65
pt 92.43 91.89 98.38 95.68 47.03 09.19 99.46 92.97 100.0 100.0 71.35 49.73 96.76 99.46 94.05 94.05 63.24 10.81
ru 98.92 98.92 097.3 94.59 02.16 02.16 94.59 91.35 96.22 97.3 23.78 14.05 98.38 99.46 88.65 85.95 08.11 07.03
et 98.92 98.92 97.84 94.05 69.73 28.65 98.38 89.73 100.0 99.46 31.89 04.32 99.46 98.92 94.59 94.05 74.05 31.89
it 95.14 95.68 95.14 97.3 63.24 16.22 99.46 98.38 100.0 100.0 58.92 42.16 99.46 99.46 94.59 93.51 65.95 07.03
ta 100.0 99.46 01.08 01.08 00.0 00.0 62.7 10.81 18.92 24.86 00.0 00.0 95.68 97.84 00.54 00.54 00.0 00.0
th 100.0 100.0 64.86 82.16 01.08 00.54 91.89 65.41 48.11 45.95 03.78 03.24 100.0 100.0 29.73 27.03 00.54 00.54
tr 99.46 100.0 68.11 67.57 30.81 017.3 87.03 85.41 99.46 097.3 47.03 30.81 96.76 98.38 71.89 73.51 29.73 15.68
Table 18: Language Steering Shift (LSS) scores from zh to languages listed for different steering factors across models. Bolded texts denote the highest values within a model.
Factor 1 Factor 2 $\Delta$ Mean p-value Sig.
pmax pmed -0.0109 0.9813 No
pmax =max -0.1789 0.0000 Yes
pmax +max -0.1946 0.0000 Yes
pmed =max -0.1680 0.0000 Yes
pmed +max -0.1837 0.0000 Yes
=max +max -0.0157 0.9117 No
Table 19: Tukey HSD pairwise comparisons between LSS scores of steering factors on LAPE neurons. Sig. denotes the significance of Factor2 compared to Factor1.
Lang SeaLLMs3-1.5B Qwen2.5-0.5B Gemma2-2B
et 0.87 $\to$ 1.50 0.49 $\to$ 0.68 0.45 $\to$ 0.46
fr 1.43 $\to$ 18.49 0.52 $\to$ 12.94 0.76 $\to$ 2.96
ht 1.01 $\to$ 0.12 0.55 $\to$ 0.89 0.42 $\to$ 0.45
id 1.92 $\to$ 15.95 0.65 $\to$ 5.52 0.55 $\to$ 0.55
it 1.15 $\to$ 10.53 0.63 $\to$ 1.26 0.48 $\to$ 1.69
jp 0.11 $\to$ 0.26 0.06 $\to$ 0.15 0.02 $\to$ 0.02
ms 1.54 $\to$ 9.70 0.61 $\to$ 3.83 0.49 $\to$ 0.60
nl 1.20 $\to$ 2.25 0.55 $\to$ 2.79 0.45 $\to$ 0.62
pt 1.47 $\to$ 15.57 0.61 $\to$ 8.36 0.61 $\to$ 2.42
qu 0.93 $\to$ 0.24 0.46 $\to$ 0.76 0.44 $\to$ 0.46
ru 0.75 $\to$ 10.06 5.22 $\to$ 4.76 0.38 $\to$ 0.48
sw 1.07 $\to$ 0.17 0.51 $\to$ 0.89 0.45 $\to$ 0.46
ta 0.64 $\to$ 0.51 0.39 $\to$ 0.18 0.40 $\to$ 0.38
th 0.56 $\to$ 2.02 0.25 $\to$ 1.44 0.17 $\to$ 0.23
tr 1.02 $\to$ 3.44 0.48 $\to$ 1.62 0.45 $\to$ 0.44
vi 1.65 $\to$ 18.38 0.52 $\to$ 10.72 0.33 $\to$ 0.66
zh 0.32 $\to$ 0.33 0.20 $\to$ 0.22 0.09 $\to$ 0.11
Table 20: BLEU score changes of self-language intervention with pmedian from en to target listed on FLORES-200 dataset. The translation prompts are non-targeted.
Lang Qwen2.5-7B Qwen2.5-0.5B SeaLLMs3-7B SeaLLMsv3-1.5B Gemma2-2B Gemma2-9b
et 0.76 $\to$ 2.54 0.49 $\to$ 0.89 0.43 $\to$ 2.99 0.87 $\to$ 1.87 0.45 $\to$ 0.52 0.24 $\to$ 0.23
fr 13.75 $\to$ 28.15 0.52 $\to$ 12.90 0.32 $\to$ 25.01 1.43 $\to$ 19.82 0.76 $\to$ 9.08 0.57 $\to$ 4.54
ht 0.92 $\to$ 1.68 0.55 $\to$ 1.02 0.34 $\to$ 1.35 1.01 $\to$ 0.89 0.42 $\to$ 0.50 0.24 $\to$ 0.26
id 1.13 $\to$ 25.01 0.65 $\to$ 6.37 0.52 $\to$ 26.88 1.92 $\to$ 15.99 0.55 $\to$ 1.64 0.29 $\to$ 0.50
it 1.16 $\to$ 14.60 0.63 $\to$ 2.37 0.44 $\to$ 15.99 1.15 $\to$ 10.77 0.48 $\to$ 4.14 0.25 $\to$ 0.45
jp 0.09 $\to$ 0.15 0.06 $\to$ 0.15 0.08 $\to$ 0.14 0.11 $\to$ 0.35 0.02 $\to$ 0.11 0.03 $\to$ 0.01
ms 1.01 $\to$ 0.63 0.61 $\to$ 3.93 0.50 $\to$ 19.14 1.54 $\to$ 10.59 0.49 $\to$ 1.18 0.31 $\to$ 0.39
nl 0.85 $\to$ 10.69 0.55 $\to$ 2.86 0.41 $\to$ 11.42 1.20 $\to$ 6.56 0.45 $\to$ 4.95 0.24 $\to$ 0.30
pt 1.33 $\to$ 25.29 0.61 $\to$ 10.06 0.58 $\to$ 19.05 1.47 $\to$ 16.80 0.61 $\to$ 10.72 0.44 $\to$ 12.90
qu 0.75 $\to$ 1.12 0.46 $\to$ 0.76 0.63 $\to$ 0.74 0.93 $\to$ 0.92 0.44 $\to$ 0.56 0.34 $\to$ 0.50
ru 2.20 $\to$ 16.23 5.22 $\to$ 4.28 4.77 $\to$ 16.36 0.75 $\to$ 10.37 0.38 $\to$ 0.35 0.24 $\to$ 1.85
sw 0.78 $\to$ 1.79 0.51 $\to$ 0.86 0.36 $\to$ 1.95 1.07 $\to$ 1.02 0.45 $\to$ 0.48 0.26 $\to$ 0.20
ta 0.45 $\to$ 0.22 0.39 $\to$ 0.22 0.29 $\to$ 0.41 0.64 $\to$ 0.43 0.40 $\to$ 0.39 0.23 $\to$ 0.13
th 0.35 $\to$ 2.53 0.25 $\to$ 1.03 0.43 $\to$ 2.96 0.56 $\to$ 2.09 0.17 $\to$ 0.18 0.13 $\to$ 0.12
tr 0.77 $\to$ 7.73 0.48 $\to$ 1.19 0.43 $\to$ 5.72 1.02 $\to$ 3.38 0.45 $\to$ 0.51 0.28 $\to$ 0.34
vi 0.74 $\to$ 26.00 0.52 $\to$ 10.11 0.36 $\to$ 21.76 1.65 $\to$ 17.31 0.33 $\to$ 1.10 0.16 $\to$ 4.57
zh 0.19 $\to$ 0.47 0.20 $\to$ 0.17 0.24 $\to$ 1.18 0.32 $\to$ 0.22 0.09 $\to$ 0.09 0.09 $\to$ 0.12
Table 21: BLEU score changes of self-language intervention from en to targets listed on FLORES-200 dataset. The translation prompts are non-targeted.
Lang SeaLLMs3-1.5B Qwen2.5-0.5B Gemma2-2B
et 2.86 $\to$ 0.35 1.46 $\to$ 0.18 0.45 $\to$ 0.01
fr 23.58 $\to$ 0.16 14.02 $\to$ 0.13 0.76 $\to$ 0.15
ht 1.28 $\to$ 0.12 0.82 $\to$ 0.09 0.42 $\to$ 0.13
id 22.06 $\to$ 0.14 7.93 $\to$ 0.03 0.55 $\to$ 0.14
it 12.44 $\to$ 0.18 6.51 $\to$ 0.09 0.48 $\to$ 0.09
jp 0.13 $\to$ 0.00 0.07 $\to$ 0.01 0.02 $\to$ 0.00
ms 17.05 $\to$ 0.13 5.04 $\to$ 0.08 0.49 $\to$ 0.14
nl 10.72 $\to$ 0.23 4.77 $\to$ 0.12 0.45 $\to$ 0.12
pt 21.99 $\to$ 0.24 13.76 $\to$ 0.07 0.61 $\to$ 0.14
qu 1.18 $\to$ 0.01 0.39 $\to$ 0.05 0.44 $\to$ 0.06
ru 13.48 $\to$ 0.32 6.96 $\to$ 0.09 0.38 $\to$ 0.17
sw 1.67 $\to$ 0.14 0.61 $\to$ 0.00 0.45 $\to$ 0.06
ta 0.33 $\to$ 0.00 0.12 $\to$ 0.00 0.40 $\to$ 0.02
th 3.59 $\to$ 0.00 1.54 $\to$ 0.03 0.17 $\to$ 0.00
tr 4.56 $\to$ 0.15 2.21 $\to$ 0.05 0.45 $\to$ 0.14
vi 22.37 $\to$ 1.42 12.74 $\to$ 0.06 0.33 $\to$ 0.14
zh 0.66 $\to$ 0.00 0.12 $\to$ 0.09 0.09 $\to$ 0.02
Table 22: BLEU score changes of self-language pmax intervention on Baseline neurons from en to target listed on FLORES-200 dataset. The translation prompts are targeted.
Figure 9: BLEU deltas for targeted pmax intervention in Qwen2.5 0.5B. Row $i$ is the target language and column $j$ is the intervention language. For example, row $x$ col $y$ represents BLEU score when translating English to language $x$ intervened by $y$ neurons.
Lang Gemma2 SeaLLMs3 Qwen2.5
9B 2B 7B 1.5B 7B 0.5B
en 70.58 68.34 86.88 82.67 85.51 75.96
fr 65.06 56.63 66.27 77.11 72.29 61.45
jp 66.11 59.02 68.82 64.44 72.99 57.14
pt 66.54 61.60 77.57 68.44 79.85 62.36
ru 71.11 65.08 73.02 67.62 70.16 58.10
zh 70.24 66.27 80.95 76.39 81.75 64.88
Table 23: Baseline performance of XWinograd across models.
Lang Gemma2 SeaLLMsv3 Qwen2.5
9B 2B 7B 1.5B 7B 0.5B
et 57.8 49.4 50.8 51.2 50.4 48.8
ht 52.4 46.8 50.4 51.4 51.8 51.2
id 69.4 62.0 73.6 65.4 73.4 57.2
it 69.6 63.6 72.0 62.0 76.6 52.2
qu 51.0 50.2 50.0 51.2 49.0 50.4
sw 61.8 54.0 53.4 53.4 53.4 54.4
ta 57.8 54.2 57.6 53.2 56.2 56.6
th 56.8 53.0 59.0 57.4 61.6 55.6
tr 64.2 54.4 61.2 54.8 61.4 53.6
vi 67.6 59.2 74.4 65.8 78.0 58.2
zh 72.6 67.2 78.0 70.0 79.8 63.6
Table 24: Baseline performance of XCOPA across models.
Lang Gemma2 SeaLLMsv3 Qwen2.5
9B 2B 7B 1.5B 7B 0.5B
et 69.4 33.3 38.8 23.5 42.1 32.8
fr 66.4 47.6 53.6 40.0 57.6 41.6
id 65.3 51.6 57.3 50.4 59.7 41.1
it 70.9 54.6 61.8 51.8 72.9 41.4
jp 71.9 53.8 63.9 55.0 69.1 48.2
ms 59.4 44.6 49.0 38.6 53.0 33.7
nl 72.1 54.2 60.6 44.6 64.1 32.7
pt 63.6 47.8 56.9 42.3 61.7 38.3
ru 55.2 45.6 52.8 42.5 55.6 40.1
ta 45.6 32.8 30.0 33.6 32.4 28.0
tr 57.0 39.8 45.8 31.7 51.0 30.5
vi 62.4 46.0 58.0 43.6 56.8 40.8
zh 55.9 45.7 73.5 73.1 77.9 55.5
Table 25: Baseline performance of Include-lite across models.
Factor Mean LSS
pmax 0.9204
pmedian 0.9095
=max 0.7415
+max 0.7258
=0 0.4109
=10p 0.1640
Table 26: Mean LSS throughout all languages and models for each steering factor.
Lang Gemma2 SeaLLMs3 Qwen2.5
9B 2B 7B 1.5B 7B 0.5B
en 9 65 49 97 30 114
nl 61 223 352 672 126 238
id 43 126 108 382 56 137
ms 48 178 131 377 71 203
vi 52 214 216 435 86 224
jp 139 495 351 690 309 573
zh 72 362 299 887 196 436
fr 42 204 343 897 109 220
pt 33 189 354 633 75 195
ru 64 262 226 444 135 256
et 156 638 176 504 355 806
ht 71 968 73 383 210 832
it 45 185 327 637 107 203
qu 172 1644 110 434 333 1426
sw 177 1231 184 551 435 1297
ta 550 1127 325 457 1085 1053
th 155 311 267 507 164 292
tr 132 286 159 668 219 355
Table 27: Proportions of LAPE neurons obtained from FLORES-200 dataset.
Lang Gemma2 SeaLLMs3 Qwen2.5
9B 2B 7B 1.5B 7B 0.5B
en 6 3995 8450 13838 2759 3653
nl 32 7051 10085 17462 5214 6570
id 22 6730 9724 16126 4864 6318
ms 27 7393 10188 17815 5324 6847
vi 19 6271 10828 20036 4303 5878
jp 62 7119 11032 19553 4996 6624
zh 38 4268 10246 19029 2935 3799
fr 29 6592 10608 17814 4683 6153
pt 21 6407 10692 17069 4547 6036
ru 32 5903 10554 17496 4231 5420
et 58 12607 11542 19546 7718 11544
ht 16 12497 11896 21920 6583 11363
it 28 7173 10945 17064 5352 6615
qu 35 14061 13129 25239 7579 12479
sw 47 14225 11856 19813 8511 13050
ta 120 21817 14465 27403 19088 19088
th 33 8352 12832 24646 5573 7692
tr 68 9139 10740 19227 6905 8502
Table 28: Proportions of baseline neurons obtained from FLORES-200 dataset.
Figure 10: Symmetric heatmaps of overlap neuron proportions in SeaLLMs3 1.5B using jaccard distance for LAPE neurons.
Figure 11: Symmetric heatmaps of overlap neuron proportions in Gemma2 2B using jaccard distance for LAPE neurons.
Figure 12: Symmetric heatmaps of overlap neuron proportions in Gemma2 9B using jaccard distance for LAPE neurons.
Figure 13: Symmetric heatmaps of overlap neuron proportions in Qwen2.5 0.5B using jaccard distance for LAPE neurons.
Figure 14: Symmetric heatmaps of overlap neuron proportions in Qwen2.5 7B using jaccard distance for LAPE neurons.
Figure 15: Symmetric heatmaps of overlap neuron proportions in SeaLLMs3 1.5B using jaccard distance for Baseline neurons.
Figure 16: Symmetric heatmaps of overlap neuron proportions in Gemma2 2B using jaccard distance for Baseline neurons.
Figure 17: Symmetric heatmaps of overlap neuron proportions in Gemma2 9B using jaccard distance for Baseline neurons.
Figure 18: Symmetric heatmaps of overlap neuron proportions in Qwen2.5 0.5B using jaccard distance for Baseline neurons.
Figure 19: Symmetric heatmaps of overlap neuron proportions in Qwen2.5 7B using jaccard distance for Baseline neurons.
Figure 20: Delta XWinograd accuracy after steering LAPE neurons (obtained from the XWinograd dataset) with pmax for Qwen2.5 0.5B.
(a) SeaLLMv3 1.5B
(b) SeaLLMv3 7B
Figure 21: Delta XWinograd accuracy after steering LAPE neurons with pmax for SeaLLMv3 1.5B and SeaLLMv3 7B. Row i is the initial language, column j is the language of intervention.
(a) Qwen2.5 0.5B
(b) Qwen2.5 7B
Figure 22: Delta XWinograd accuracy after steering LAPE neurons with pmax for Qwen2.5 0.5B and Qwen2.5 7B. Row i is the initial language, column j is the language of intervention.
(a) Gemma2 2B
(b) Gemma2 9B
Figure 23: Delta XWinograd accuracy after steering LAPE neurons with pmax for Gemma2 2B and 9B. Row i is the initial language, column j is the language of intervention.
Figure 24: Delta XWinograd accuracy after steering LAPE neurons with pmedian for Gemma2 2B. Row i is the initial language, column j is the language of intervention.
Figure 25: Delta XWinograd accuracy after steering LAPE neurons with pmedian for Qwen2.5 0.5B. Row i is the initial language, column j is the language of intervention.
Figure 26: Delta XWinograd accuracy after steering LAPE neurons with pmedian for SeaLLMs 1.5B. Row i is the initial language, column j is the language of intervention.
Figure 27: Delta XWinograd accuracy after steering LAPE neurons with pmedian for Gemma2 9B. Row i is the initial language, column j is the language of intervention.
Figure 28: Delta XWinograd accuracy after steering LAPE neurons with pmedian for Qwen2.5 7B. Row i is the initial language, column j is the language of intervention.
Figure 29: Delta XWinograd accuracy after steering baseline neurons with pmax for SeaLLMv3 1.5B. Row i is the initial language, column j is the language of intervention.
Figure 30: Delta XWinograd accuracy after steering baseline neurons with pmax for Qwen2.5 0.5B. Row i is the initial language, column j is the language of intervention.
Figure 31: Delta XWinograd accuracy after steering baseline neurons with pmax for Gemma2 2B. Row i is the initial language, column j is the language of intervention.
Figure 32: Delta XCOPA accuracy after steering LAPE neurons with pmax for Gemma2 9B.
Figure 33: Delta XCOPA accuracy after steering LAPE neurons with pmax for Gemma2 2B.
Figure 34: Delta XCOPA accuracy after steering LAPE neurons with pmax for Qwen2.5 0.5B.
Figure 35: Delta XCOPA accuracy after steering LAPE neurons with pmax for Qwen2.5 7B.
Figure 36: Delta XCOPA accuracy after steering LAPE neurons with pmax for SeaLLM3 1.5B.
Figure 37: Delta XCOPA accuracy after steering LAPE neurons with pmax for SeaLLM3 7B.
Figure 38: Delta XCOPA accuracy after steering LAPE neurons with pmedian for SeaLLMs 1.5B. Row i is the initial language, column j is the language of intervention.
Figure 39: Delta XCOPA accuracy after steering LAPE neurons with pmedian for Gemma2 2B. Row i is the initial language, column j is the language of intervention.
Figure 40: Delta XCOPA accuracy after steering LAPE neurons with pmedian for Qwen2.5 0.5B. Row i is the initial language, column j is the language of intervention.
Figure 41: Delta XCOPA accuracy after steering baseline neurons with pmax for Gemma2 2B.
Figure 42: Delta XCOPA accuracy after steering baseline neurons with pmax for Qwen2.5 0.5B.
Figure 43: Delta XCOPA accuracy after steering baseline neurons with pmax for SeaLLM3 1.5B.
Figure 44: Delta Include-lite accuracy after steering LAPE neurons with pmax for Qwen2.5 0.5B.
Figure 45: Delta Include-lite accuracy after steering LAPE neurons with pmax for Qwen2.5 7B.
Figure 46: Delta Include-lite accuracy after steering LAPE neurons with pmax for Gemma2 2B.
Figure 47: Delta Include-lite accuracy after steering LAPE neurons with pmax for Gemma2 9B.
Figure 48: Delta Include-lite accuracy after steering LAPE neurons with pmax for SeaLLMv3 1.5B.
Figure 49: Delta Include-lite accuracy after steering LAPE neurons with pmax for SeaLLMv3 7B.
Figure 50: Delta Include-lite accuracy after steering LAPE neurons with pmedian for Qwen2.5 0.5B.
Figure 51: Delta Include-lite accuracy after steering LAPE neurons with pmedian for Gemma2 2B.
Figure 52: Delta Include-lite accuracy after steering LAPE neurons with pmedian for SeaLLMv3 1.5B.
Figure 53: Delta Include-lite accuracy after steering Baseline neurons with pmax for Qwen2.5 0.5B.
Figure 54: Delta Include-lite accuracy after steering Baseline neurons with pmax for Gemma2 2B.
Figure 55: Delta Include-lite accuracy after steering Baseline neurons with pmax for SeaLLmv3 1.5B.
Figure 56: PPL changes after steering LAPE neurons with pmax for Qwen2.5 0.5B.
Figure 57: PPL changes after steering LAPE neurons with pmax for Qwen2.5 7B.
Figure 58: PPL changes after steering LAPE neurons with pmax for Gemma2 2B.
Figure 59: PPL changes after steering LAPE neurons with pmax for Gemma2 9.B
Figure 60: PPL changes after steering LAPE neurons with pmax for SeaLLM3 1.5B.
Figure 61: PPL changes after steering LAPE neurons with pmax for SeaLLM3 7B.
Figure 62: PPL changes after steering LAPE neurons with pmedian for Qwen2.5 0.5 B.
Figure 63: PPL changes after steering LAPE neurons with pmedian for SeaLLMs3 1.5 B.
Figure 64: PPL changes after steering LAPE neurons with pmedian for Gemma2 2B.
Figure 65: PPL changes after steering Baseline neurons with pmax for Qwen2.5 0.5B.
Figure 66: PPL changes after steering Baseline neurons with pmax for Gemma2 2B.
Figure 67: PPL changes after steering baseline neurons with pmax for SeaLLM3 1.5B.
Figure 68: Layer-wise average LSS scores across all languages for Gemma2 2B.
Figure 69: Layer-wise average LSS scores when intervened by en neurons for Gemma2 2B.
Figure 70: Layer-wise average LSS scores when intervened by ms neurons for Gemma2 2B.
Figure 71: Layer-wise average LSS scores when intervened by et neurons for Gemma2 2B.
Figure 72: Layer-wise average LSS scores when intervened by nl neurons for Gemma2 2B.
Figure 73: Layer-wise average LSS scores when intervened by id neurons for Gemma2 2B.
Figure 74: Layer-wise average LSS scores when intervened by vi neurons for Gemma2 2B.
Figure 75: Layer-wise average LSS scores when intervened by jp neurons for Gemma2 2B.
Figure 76: Layer-wise average LSS scores when intervened by zh neurons for Gemma2 2B.
Figure 77: Layer-wise average LSS scores when intervened by fr neurons for Gemma2 2B.
Figure 78: Layer-wise average LSS scores when intervened by pt neurons for Gemma2 2B.
Figure 79: Layer-wise average LSS scores when intervened by ru neurons for Gemma2 2B.
Figure 80: Layer-wise average LSS scores when intervened by it neurons for Gemma2 2B.
Figure 81: Layer-wise average LSS scores when intervened by ta neurons for Gemma2 2B.
Figure 82: Layer-wise average LSS scores when intervened by th neurons for Gemma2 2B.
Figure 83: Layer-wise average LSS scores when intervened by tr neurons for Gemma2 2B.
Figure 84: Visualization of token log-prob for base and target answers. Base answers are in the prompt language (en), target answers are in the language of intervention (nl, id,ms,vi,jp,zh,fr,pt).
Figure 85: Visualization of token log-prob for base and target answers in intervened and non-intervened LLMs. Base answers are in the prompt language (en), target answers are in the language of intervention (ru, et,it,ta,th,tr).
Figure 86: A sample of Qwen2.5 0.5B output for targeted prompt translation of en to nl, id,ms,fr,qu,ht, sw.
Figure 87: A sample of Qwen2.5 0.5B output for non-targeted prompt translation of en to nl, id,ms,fr,qu,ht, sw.